### v1

# From Source to Target: Leveraging Transfer Learning for Predictive Process Monitoring in Organizations

From Source to Target: Leveraging Transfer Learning for Predictive Process Monitoring in Organizations
Sven Weinzierl1∗, Sandra Zilker1,2, Annina Liessmann1, Martin Käppel1,
Weixin Wang1, and Martin Matzner1
1Chair of Digital Industrial Service Systems, Friedrich-Alexander-Universität Erlangen-Nürnberg,
Fürther Str. 248, 90429 Nürnberg, Germany
2Professorship for Business Analytics, Technische Hochschule Nürnberg Georg Simon Ohm
Hohfederstr. 40, 90489 Nürnberg, Germany
Corresponding author: Sven Weinzierl, E-mail: sven.weinzierl@fau.de
Event logs reflect the behavior of business processes that are mapped in organizational information systems. Predictive process monitoring (PPM) transforms these data into value by creating process-related predictions that provide the insights required for proactive interventions at process runtime. Existing PPM techniques require sufficient amounts of event data or other relevant resources that might not be readily available, preventing some organizations from utilizing PPM. The transfer learning-based PPM technique presented in this paper allows organizations without suitable event data or other relevant resources to implement PPM for effective decision support. The technique is instantiated in two real-life use cases, based on which numerical experiments are performed using event logs for IT service management processes in an intra- and inter-organizational setting. The results of the experiments suggest that knowledge of one business process can be transferred to a similar business process in the same or a different organization to enable effective PPM in the target context. With the proposed technique, organizations can benefit from transfer learning in an intra- and inter-organizational setting, where resources like pre-trained models are transferred within and across organizational boundaries.
Keywords: Predictive process monitoring, Process mining, Transfer learning, Language model
1 Introduction
Business processes are the backbone of organizational value creation (Dumas et al. 2018). With ongoing digitalization, business processes are increasingly mapped in information systems (IS), such as enterprise resource planning (ERP) systems, and executed by process users and other stakeholders (Beverungen et al. 2021). Thus, business processes leave large amounts of digital event data reflecting their behavior (van der Aalst 2016). To transform these data into value, various business process analytics approaches have been developed (zur Muehlen and Shapiro 2009).
One of these approaches is predictive process monitoring (PPM) – a set of techniques that leverage historical event log data (Maggi et al. 2014) to create prediction models Grigori et al. (2004), capable of predicting various targets in running business process instances, such as the next activity, remaining time, or process-related outcomes (Di Francescomarino et al. 2018). By generating these process-related predictions, PPM enables proactive interventions in running business process instances, helping to reduce risks and optimize resource allocation (Di Francescomarino et al. 2018). This leads to improved business process performance (Marquez-Chamorro et al. 2017), even across organizational boundaries (Oberdorf et al. 2023). Therefore, PPM is gaining momentum in business process management (BPM), while its techniques are considered a new class of process mining techniques (Di Francescomarino and Ghidini 2022).
Most PPM techniques rely on shallow machine learning (ML) models (Marquez-Chamorro et al. 2017) or deep learning (DL) models (Rama-Maneiro et al. 2023) to generalize knowledge from historical to unseen event log data (Weinzierl et al. 2024). These models are trained on historical event log data and then applied to new instances of a running business process in order to predict various targets depending on the business context and analytical goals (Maggi et al. 2014). However, organizations – or different departments or subsidiaries within the same organization – often lack the prerequisites for building effective prediction models. These include access to sufficient amounts of event log data (Käppel et al. 2021) and the technical infrastructure required to train complex models, such as graphical processing units.
Consequently, some organizations cannot benefit from PPM, while others may already have access to richer event log data or (pre-)trained prediction models for similar business processes. If these processes are sufficiently similar in terms of semantics (e.g., meaning of activity names) and structure (e.g., similar control flow), it becomes possible to overcome the aforementioned obstacles by transferring a prediction model from one business process to another instead of training a new model from scratch. This idea is supported by the fact that many organizations implement standardized or partially standardized business processes (e.g., information technology service management (ITSM) processes), which are mapped in ERP systems such as SAP ERP, Oracle ERP Cloud, or Microsoft Dynamics 365.
The field of transfer learning (TL) provides techniques specifically designed to support such scenarios (Pan and Yang 2009), enabling the transfer of ML models and the knowledge they contain from one learning problem (i.e., domain and task) to another. In PPM, TL can be applied in inter-organizational scenarios, in which prediction models are transferred between processes of different organizations, and in intra-organizational scenarios, where a prediction model is transferred from one process to another within a single organization (e.g., across departments or subsidiaries). In both scenarios, we consider all resources to be transferred as process knowledge. This includes the prediction model (also called the base model). Process knowledge originates in its source context, which includes the business process itself, the event data generated by it, and the specific PPM task being addressed using the event data of the business process. The transferred process knowledge is then applied in the target context (see Figure 1).
Figure 1: Transfer process knowledge from source to target contexts in PPM.
Building on this idea, we present a TL-based technique for PPM that relies on existing similarities between business processes and aims to apply a prediction model to a target context. Therefore, the prediction model is transferred together with further resources, such as information on the corresponding encoding strategy. In particular, this work contributes to the existing body of knowledge in four ways:
  1. 1.
    We present a TL-based technique for PPM that enables the transfer of a base model for an outcome-oriented prediction target without the necessity to fine-tune it on the event data of the target context. Our approach also provides guidance on which resources need to be transferred, including pre-trained models, information on the corresponding encoding strategy, further preprocessing details, and metrics used for measuring the prediction performance.
  2. 2.
    We propose a cross-domain encoding strategy for event data that enables effective knowledge transfer from source to target contexts. While timestamp information is encoded via a relative mapping approach, activity information is encoded using pre-trained embedding models to preserve semantic consistency across contexts.
  3. 3.
    We propose a two-layer long short-term memory (LSTM) model with dedicated parameter initialization as the prediction model of the source context, used for mapping the encoded traces onto process-related outcome predictions. This model improves learning from the event data of the source context and facilitates knowledge transfer to the target context.
  4. 4.
    We empirically evaluate our TL-based PPM technique in both an intra- and inter-organizational ITSM use case. The results show that the transferred base model achieves a clearly higher prediction performance than models built with traditional PPM techniques on the event data of the target context.
To carry out our research, we apply computational design science research (DSR) (Rai et al. 2017) and structure our work according to the DSR publication schema by Gregor and Hevner (2013): First, we present relevant background on ML and DL in PPM, and TL in general (Section 2). Second, we explain the practical relevance of TL for PPM in organizations and introduce inter- and intra-organizational TL in the context of PPM (Section 3). Subsequently, we give an overview of related work on TL in PPM and position our proposed technique with regard to them (Section 4). Then, we provide a detailed description of the proposed artifact (Section 5). Afterward, the evaluation of our artifact is described by introducing the intra- and inter-organizational use cases, and providing details on the setup used in the evaluation (Section 6). Then, we present the results of our evaluation (Section 7), discuss implications and future research directions (Section 8), and conclude the paper (Section 9).
2 Background
2.1 Event log data and predictive process monitoring
A business process is defined as a sequence of interrelated activities and decisions undertaken to achieve a valuable outcome for a customer (Dumas et al. 2018). Each execution of a business process is referred to as a process instance or case (van der Aalst 2016). Information technology (IT) systems (e.g., ERP systems) commonly record these executions in the form of event logs, that is, collections of timestamped events described by various event attributes that store information about the execution of activities (i.e., steps in the process) (van der Aalst 2016, p. 130). Each event contains at least a case identifier, the name of the executed activity, and a timestamp indicating when the event occurred. Events belonging to the same process instance can be temporally ordered by their timestamp to form a so-called trace (van der Aalst 2016, p. 134).
Event logs serve as the main input for PPM techniques. PPM provides a set of techniques that utilize event log data to make predictions on the future behavior or properties of ongoing process instances (Verenich et al. 2019). These techniques aim to support proactive decision-making to enhance process performance and counteract potential risks during process execution (Marquez-Chamorro et al. 2017). PPM techniques can target various types of predictions. One common prediction task is outcome prediction, which aims to forecast whether a process instance will end with a certain (often categorical) outcome (Teinemaa et al. 2019). Other prediction targets include the next activity or the remaining sequence of activities (suffix), as well as temporal aspects such as the timestamp of the next activity or the remaining time until process completion (Verenich et al. 2019).
2.2 Machine learning in predictive process monitoring
A branch of early PPM approaches creates prediction models by augmenting discovered process models (or structurally similar representations such as Petri nets or finite automatons) with predictive capabilities (e.g., van der Aalst et al. 2011; Rogge-Solti and Weske 2015). However, as real-world processes are usually more complex than reflected in discovered process models, which frequently fail to capture all execution scenarios, their predictive power is limited (van der Aalst 2011). To overcome this limitation, a more recent branch of PPM approaches emerged that exploit the generalization capabilities of ML in identifying and generalizing patterns in event log data to learn prediction models in a data-driven manner.
Besides a wide range of conventional ML algorithms such as linear regressions, logistic regressions, decision trees, random forests, and gradient boosted decision trees, deep neural networks (DNNs), from the field of DL, have been applied for building prediction models (Marquez-Chamorro et al. 2017; Teinemaa et al. 2019; Verenich et al. 2019). Especially, the latter has gained prominence due to its ability to learn complex patterns and dependencies in large datasets (LeCun et al. 2015). DL techniques have shown superior prediction performance and generalizability across various PPM tasks (Kratsch et al. 2021), and have thus become the state of the art.
In recent years, a variety of neural network architectures have been employed for different prediction targets, including multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), and recurrent neural network (RNN) with either LSTM cells or gated recurrent units (GRUs), with LSTM networks being favored by the majority (Neu et al. 2022). Among these, RNN, particularly LSTM networks, have been the most widely adopted in PPM. For example, Evermann et al. (2017) applied LSTMs for next activity prediction, while Tax et al. (2017) demonstrated that LSTMs outperform conventional ML approaches in predicting both the next activity and its corresponding timestamp. More recently, transformer-based models, originally introduced in natural language processing, have shown high prediction performance. These models utilize self-attention mechanisms to effectively capture long-range dependencies in sequential data, as present in event data. For instance, Zaharah A. Bukhsh (2021) proposed a transformer model capable of predicting next activities, remaining time, and timestamps. Beyond efforts to enhance prediction performance through increasingly sophisticated and powerful neural network architectures, complementary strategies have focused on enriching the input data by incorporating additional sources of information. While many prediction models primarily rely on control-flow information and time information derived from timestamps (e.g., elapsed time), several approaches extend this by integrating contextual information, for example, in the form of case-level or resource attributes (Weinzierl et al. 2019). For example, Di Francescomarino et al. (2018) proposed multi-perspective models that combine control flow and contextual information for outcome prediction. As another example, Hinkka et al. (2019) suggested clustering categorical event attributes and using the resulting cluster labels as additional information in a RNN. In contrast to these approaches, Pegoraro et al. (2021) and Cabrera et al. (2022) integrate unstructured data in the form of text as contextual information. The ability of DL methods to handle diverse prediction tasks and to incorporate heterogeneous data sources not only demonstrates their flexibility but also hints at the potential for reusing learned representations in varying contexts (Weinzierl et al. 2024).
To build a prediction model using an ML or DL algorithm, it is crucial to appropriately encode the underlying event data, as the quality and structure of the encoding strongly affect the model’s ability to generalize. Thus, the encoding determines how transferable and reusable the learned representations are across different but related tasks. In many approaches, process instances are transformed into fixed-size feature vectors labeled with a prediction target, as required in supervised ML settings such as classification or regression (Di Francescomarino and Ghidini 2022). However, especially in DL-based models, alternative representations such as event sequences or graph-structured inputs are also employed, allowing models to process richer structural information directly (Weinzierl 2022). Typically, two essential encoding steps are necessary: first, sequence encoding, to capture the sequential structure of the traces, and second, event encoding, to encode the information contained in the events (i.e., the event attributes). For the sequence encoding, various strategies have been proposed. These strategies range from simple boolean encodings indicating the presence of specific activities to more detailed representations that preserve the full control flow by including the complete sequence of events or the last $n$ events (states) (Di Francescomarino and Ghidini 2022). The encoding of the event attributes depends largely on their data types. Categorical attributes, such as activity names or resources, are either one-hot encoded or represented by an embedding vector, that is, a dense vector capturing the semantics. Numerical attributes usually require only normalization or scaling into a particular interval (Teinemaa et al. 2019). Temporal information, especially timestamp attributes, can be encoded by deriving features, such as time since the last event or the weekday, which then can be treated as numerical or categorical values (Tax et al. 2017). Additional data attributes on the process context can be present in the form of text. For these attributes, simple natural language processing techniques can be used, such as bag-of-n-grams or latent Dirichlet allocation (Teinemaa et al. 2016), to extract semantically shallow features from text data. By using more advanced techniques that are based on (large) language models, such as contextualized word embedding models like bidirectional encoder representations from transformers (BERT) or robustly optimized BERT pretraining approach (RoBERTa), semantically rich features can be extracted from text data Cabrera et al. (2022); Liessmann et al. (2024).
2.3 Transfer learning
Typical ML-based approaches, like the ML-based PPM techniques described in Section 2.2, require the availability of labeled data with the same distribution in training and test sets (Pan and Yang 2009). This might not be the case in practice. TL offers a solution, as it “allows the domains, tasks, and distributions used in training and testing to be different” (Pan and Yang 2009, p. 1346). A domain $\pazocal{D}=\{\pazocal{X},P(X)\}$ is described by the feature space $\pazocal{X}$, that is, the set of all term vectors, and the marginal probability distribution $P(X)$, where $X$ is an instance set $X=\left\{\mathbf{x}\mid\mathbf{x}_{i}\in\pazocal{X},\text{with }i=1,\dots,n\right\}$. Task $\pazocal{T}$ consists of the label space $\pazocal{Y}$, that is, the set of all labels, and the predictive function $f(\cdot)$. In TL, a prediction model for the domain with limited data, that is, the target domain, can be improved by transferring knowledge from another domain, the source domain (Weiss et al. 2016). More specifically, the learning of the target prediction function $f_{T}(\cdot)$ for the target learning task $\pazocal{T}_{T}$ and target domain $\pazocal{D}_{T}$ can be improved by transferring knowledge from the source domain $\pazocal{D}_{S}$ and source task $\pazocal{T}_{S}$ with either $\pazocal{D}_{S}\neq\pazocal{D}_{T}$ or $\pazocal{T}_{S}\neq\pazocal{T}_{T}$ (Pan and Yang 2009).
In the context of PPM, a domain is represented by the encoded event data $X$. The feature space $\pazocal{X}$ of the encoded event data $X$ is determined by the selected event attributes, the type of sequence encoding, and the type of event encoding. The marginal distribution $P(X)$ of the encoded event data $X$ is the distribution of the features used to represent the event data $X$. The addressed PPM task (e.g., the prediction of a process-related outcome) is denoted by $\pazocal{T}$. Its goal is to learn a prediction function $f(\cdot)$ that maps a running process instance to a label $y\in\pazocal{Y}$, where $y$ is a value and $\pazocal{Y}$ is the space of all possible values, respectively, for the PPM task $\pazocal{T}$.
Figure 2: Categories of TL (Zhuang et al. 2020).
The field can be categorized from a problem and a solution perspective, as shown in Figure 2 (Zhuang et al. 2020). From the problem perspective, depending on the existence of the label, three categories exist: i) transductive TL (label only present in $\pazocal{D}_{S}$, $\pazocal{D}_{S}\neq\pazocal{D}_{T}$, $\pazocal{T}_{S}=\pazocal{T}_{T}$), ii) inductive TL (label present in $\pazocal{D}_{T}$, $\pazocal{T}_{S}\neq\pazocal{T}_{T}$), and iii) unsupervised TL (label unknown in both domains, $\pazocal{T}_{S}\neq\pazocal{T}_{T}$) (Pan and Yang 2009). Another differentiation can be made by comparing the feature spaces of the source and target domains. While homogeneous TL refers to $\pazocal{X}^{S}=\pazocal{X}^{T}$, in heterogeneous TL, the feature spaces of the source and the target domains are not the same ($\pazocal{X}^{S}\neq\pazocal{X}^{T}$) (Weiss et al. 2016). From a solution perspective, Pan and Yang (2009) suggest four categories of approaches based on what is transferred: i) instance-based, ii) feature-based, iii) parameter-based, and iv) relational-based. In instance-based approaches, instances from the source domain are re-weighted to be used in the target domain. Feature-based approaches aim to transfer knowledge with a learned feature representation. This can be done asymmetrically, meaning transforming source features to align with target features, or symmetrically, which means discovering a common new feature space (Zhuang et al. 2020). In parameter-based approaches, shared (hyper-)parameters or hyperpriors between the source and target domains are transferred. Lastly, relational-based approaches transfer knowledge in the form of relations between the source and target data, assuming that the data are not independent and identically distributed (Pan and Yang 2009).
Since we aim to transfer process knowledge within and across organizational boundaries, data privacy issues might prohibit instance-based approaches. Assuming that the organizations concerned are operating independently and not identically, relational-based approaches might not be appropriate for our setting. However, feature-based and parameter-based approaches seem promising to transfer process knowledge within and across organizational boundaries.
3 Practical relevance of transfer learning for predictive process monitoring in organizations
Transferring and utilizing knowledge from a source context S to a target context T is particularly interesting in connection with business processes. In organizations, IT systems are in place to support the execution of business processes to generate value. The widespread use of ERP systems offering similar modules for purchasing, sales and distribution, production, warehouse and inventory management, human resource management, or financial and cost accounting makes business processes comparable across organizations. Alongside ERP systems, the use of function-specific software, such as Salesforce for customer relationship management or ServiceNow for IT service management, leads to a standardization of processes. In addition to IT systems that offer a common ground for the execution of business processes, organizations can follow industry best practices and standards, for example, by the International Organization for Standardization (ISO)  [^1] https://www.iso.org (Accessed 23 July 2025) or the International Electrotechnical Commission (IEC)  [^2] https://www.iec.ch/homepage (Accessed 23 July 2025). In IT, standards include ISO/IEC 27001 for information security management systems or the Information Technology Infrastructure Library (ITIL)  [^3] https://www.axelos.com/certifications/itil-service-management/ (Accessed 23 July 2025) for IT service management. In manufacturing, ISO 9001 for quality management or Six Sigma for process improvement are common. Besides the voluntary adherence to industry standards, the activities and processes of an organization are also governed by rules and regulations. Prime examples of such regulations and directives are General Data Protection Regulation (GDPR)  [^4] https://gdpr.eu (Accessed 23 July 2025) and the recently introduced Corporate Sustainability Reporting Directive (CSRD)  [^5] https://finance.ec.europa.eu (Accessed 23 July 2025). Similarly, organizations must follow certain accounting standards based on their location, such as Generally Accepted Accounting Principles (GAAP)  [^6] https://fasab.gov/accounting-standards/ (Accessed 23 July 2025) or International Financial Reporting Standards (IFRS)  [^7] https://www.ifrs.org (Accessed 23 July 2025). Lastly, organizations have similar objectives based on which their processes are designed. For example, an organization can strive for cost or time reduction, improved efficiency, increased product or service quality, or customer satisfaction in their process executions. The use of comparable software tools, adherence to regulations, and (industry) standards, as well as common objectives, make different business processes within and across organizations increasingly similar to each other.
Figure 3: Intra- and inter-organizational TL for PPM. [^9] For simplification, Figure 9 shows the transfer of the prediction model representative for all transferred resources from the source to the target context.
This offers opportunities to transfer process knowledge within one organization (see Figure 9). Intra-organizational TL for PPM allows an organization to transfer knowledge across its processes. This might be especially useful when considering an organization operating with multiple subsidiaries in various locations. Such conglomerates often maintain multiple ERP systems for their subsidiaries based on industry or location. For example, an organization C has PPM with event data from its ERP system already in use in a source subsidiary $C_{S}$. To harness existing process knowledge in subsidiary $C_{S}$, a target subsidiary $C_{T}$ can utilize subsidiary $C_{S}$’s prediction models for its process data. This becomes inherently convenient when laws and regulations inhibit the collection of sufficient historical process data for subsidiary $C_{T}$ to train its own prediction models. In addition, intra-organizational TL for PPM offers the possibility for an organization to go beyond PPM pilot projects and provide process owners with the tools to implement PPM by themselves. Pre-trained prediction models for particular processes can be accessed by process owners to apply PPM without the need for costly resources and extensive model training, almost like an internal model-as-a-service offering.
While transferring process knowledge through TL within one organization mainly benefits the same organization, TL for PPM can also be applied across organizational boundaries (see Figure 9). We coin this inter-organizational TL for PPM. For example, a target organization T is interested in the application of PPM and is planning to launch a pilot project to investigate its use for one of its support processes, incident management. Organization T is logging all information related to solving incidents in its service desk using ServiceNow, following ITIL while adhering to ISO/IEC 27001. With PPM, organization T aims to reduce the resolution time. To keep resources and costs low and to be able to launch a proof-of-concept relatively fast, organization T can use a transferred prediction model, trained with event data of a similar incident management process of source organization S instead of training a prediction model from scratch. Another use case for inter-organizational TL for PPM represents the acquisition of organization T by organization S. Some processes of organization S are already supported by PPM. As part of the merger, organization S wants to integrate organization T and its ERP system in its PPM efforts. To enable smooth integration, the prediction models used for processes in organization S can be transferred to similar processes in organization T.
The increasing similarities of business processes create an application area for intra- and inter-organizational TL in PPM. As inter-organizational TL in PPM aims to transfer process knowledge across organizations, the domains (i.e., the feature space and marginal probability distribution of event data) of the source and target contexts tend to differ more than in intra-organizational TL in PPM, in which knowledge is transferred between business processes within the same organization. This greater difference in the source and the target domains in both application cases can be attributed to, for example, organization-specific standards and regulations for performing the same type of business process.
4 Related work
Related TL-based PPM techniques can generally be categorized into two groups: i) PPM techniques that transfer resources within one business process in an organization (i.e., $\pazocal{D}_{S}$ and $\pazocal{D}_{T}$ represent event data on the same business process) and ii) PPM techniques that transfer resources between similar business processes within one organization or across different organizations (i.e., $\pazocal{D}_{S}$ and $\pazocal{D}_{T}$ represent event data on similar business processes within the same or different organizations).
Some PPM techniques in the first group aim to transfer parameters between models in transition system-based approaches, to capture patterns in event data better, and ultimately improve prediction performance. Cao et al. (2022) present an approach for remaining time prediction. This approach combines Petri net-based process modeling and DL. A Petri net is mined from event log data to construct a transition sequence, and prefixes are classified into partitions (transition division) based on their last activity. Autoencoder models are then trained on the prefixes within each partition to reduce dimensionality. Based on the prefixes encoded via the corresponding autoencoder model, an MLP model is trained for each partition. TL is utilized by transferring parameters from the MLP model trained on the event data of the previous partition ($\pazocal{D}_{S}$) to the MLP model of the current partition, and fine-tuning the MLP model of the current partition with available event data ($\pazocal{D}_{T}$). In a similar approach, Ni et al. (2022) introduce an autoencoder-based transition system for predicting the remaining time of a process. They mine a transition system from event log data and use a sparse autoencoder to compactly represent the event log traces. For each state of the transition system, an MLP is pre-trained on the encoded event data of all states ($\pazocal{D}_{S}$) to capture global dependencies. These models’ parameters are then transferred to new models, which are fine-tuned using state-specific event data ($\pazocal{D}_{T}$).
Other PPM techniques in the first group aim to transfer parameters between DL models to improve prediction performance. Mehdiyev et al. (2020) propose a technique for the next activity prediction, in which two autoencoder models are trained in a semi-supervised manner on available event data ($\pazocal{D}_{S}$). The parameters from the hidden layers of these models are then transferred to a prediction model for the next activity prediction, which is subsequently fine-tuned using the same event data ($\pazocal{D}_{T}$). The work of Folino et al. (2019) suggests a technique for outcome prediction with limited labeled event data. An LSTM-encoder model is trained on event data ($\pazocal{D}_{S}$) for next activity prediction, as no extra labels are required for this target. The last hidden state of the encoder model’s final LSTM layer is then transferred to a second model for outcome prediction. This model is then fine-tuned using the same event data ($\pazocal{D}_{T}$) for which a process-related outcome label is available.
Besides parameter-based TL, Cabrera et al. (2022) propose a PPM technique that leverages feature-based TL. This is achieved by using various BERT models – pre-trained on a huge corpus of text data ($\pazocal{D}_{S}$) – to encode the text information available in event data ($\pazocal{D}_{T}$) and adding the encoded information to the input of a multi-tasking LSTM model. This LSTM model addresses the next activity prediction and the next timestamp prediction. In addition, the work suggests that fine-tuning the pre-trained BERT models on available event data using the next activity prediction task enhances the LSTM model’s prediction performance for next activities and timestamps as it adapts the representation of the BERT models to the process control-flow.
Some PPM techniques in the first group also perform both parameter- and feature-based TL. Pfeiffer et al. (2021) propose a PPM technique combining Gramian angular fields and representation learning. The Gramian angular fields are applied to encode the event data as 2D images, while representation learning is performed to obtain a generic process representation from the encoded event data ($\pazocal{D}_{S}$). This representation is obtained using a two-step approach. First, a CNN model is trained to map the encoded input onto trace variants. Second, the pre-trained CNN model is extended with a fully-connected layer and heads for various prediction tasks, such as the next activity and next timestamp prediction, and then fine-tuned on the encoded event data. Subsequently, all heads except the desired one are removed, and the model is fine-tuned again on the encoded input ($\pazocal{D}_{T}$). By learning a model capturing a generic process representation and aligning it to various prediction tasks, feature- and parameter-based TL is performed within one model. In a similar approach, Chen et al. (2022) propose to learn a generic representation of traces by pre-training a BERT on a set of unlabeled trace data ($\pazocal{D}_{S}$) using a custom masked activity modeling task. The pre-trained BERT model is then fine-tuned on another set of labeled trace data ($\pazocal{D}_{S}$) for next activity prediction or outcome prediction. As in the approach of Pfeiffer et al. (2021), one model capturing a generic process representation is learned and aligned to prediction tasks. Therefore, this approach also performs feature- and parameter-based TL. Pasquadibisceglie et al. (2023) suggest using TL for handling concept drift in PPM. Specifically, once a drift is detected within event data ($\pazocal{D}_{S}$), the parameters of both an LSTM model for the next activity prediction and a word2vec model for activity encoding are updated with current event data ($\pazocal{D}_{T}$). This improves the prediction performance of the LSTM model and enables the word2vec to deal with newly occurring activities. However, even though these PPM techniques incorporate parameter- and feature-based TL, they are not designed to work with different business processes, reflected by general differences in event attribute values and distributions in the source and the target domains, which is necessary for intra- or inter-organizational TL settings.
The second group of PPM techniques aims to transfer resources from one business process to another. van Luijken et al. (2023) use event data $\pazocal{D}_{S}$ of a business process in the source context to pre-train an LSTM model and a generative pre-trained transformer (GPT) model for predicting activity and timestamp suffixes of a prefix. The parameters of these pre-trained models are then transferred to new models in the target context that address the same prediction tasks. After this transfer, some layers remain unchanged and others are fine-tuned using the event data $\pazocal{D}_{T}$ of a business process in the target context. Although the authors demonstrate the feasibility of heterogeneous TL with real-life event data by transferring model parameters from source to target contexts, their technique does not consider the embedding of different sets of activities in event data $\pazocal{D}_{S}$ and $\pazocal{D}_{T}$. However, feature-based TL is – besides parameter-based TL– necessary to transfer process knowledge between similar business processes within the same or across different organizations, as the values and distributions of event attributes in the event data of the business processes can differ from each other.
Table 1: Overview of existing works applying TL for PPM.
Authors Problem categorization Solution categorization Prediction model Input data Transfer goal Different business processes Activity encoding Timestamp encoding Fine-tuning in target not necessary Prediction target Data set
Folino et al. (2019) Inductive, homogeneous Parameter-based LSTM-encoder model ACT, RES, DA Transfer process knowledge from a prediction model tailored to an auxiliary task, for which no labeled event data is required, to an prediction model tailored to a target prediction task, for which labeled event data is required, but scarcely available One-hot encoding Extraction of six time-related features (hour, weekday, month, and times elapsed since the process instance’s start, previous event, and midnight), min-max normalization OP, NAP BPIC12
Pfeiffer et al. (2021) Inductive, homogeneous Parameter-based, feature-based CNN model with task-specific layer ACT, TS, RES, DA Transfer process knowledge from a model capturing a generic business process representation for various single prediction tasks to prediction models for single tasks to address them more effectively Integer encoding, min-max scaling, Gramian angular field transformation Extraction of one time-related feature (duration since start), min-max normalization, Gramian angular field transformation RTP, NAP, OP Helpdesk, BPIC12, BPIC13, BPIC17, BPIC20, MobIS
Chen et al. (2022) Inductive, homogeneous Parameter-based, feature-based Multi-tasking, BERT-based transformer model ACT Transfer process knowledge from a generic business process representation for multiple tasks to prediction models for single tasks to address them more effectively Embedding layer, positional encoding - NAP, OP Helpdesk, BPIC12, Sepsis
Ni et al. (2022) Transductive, homogeneous Parameter-based MLP model ACT, TS, RES, DA Transfer process knowledge from a generic prediction model pre-trained on event data (all states in a transition system) to a specific prediction model fine-tuned on event data (one state), to address a prediction task more effectively One-hot encoding, autoencoder Extraction of four time-related features (year, month, week, and day), discretization, autoencoder RTP Helpdesk, Hospital Billing, BPIC12, BPIC13
Cao et al. (2022) Transductive, homogeneous Parameter-based MLP model ACT, TS Transfer process knowledge from a prediction model pre-trained on event data (previous transition division) to another prediction model fine-tuned on event data (current transition division), to address a prediction task more effectively One-hot encoding, autoencoder Extraction of eight time-related features (e.g., month, execution time, elapsed time), min-max normalization, autoencoder RTP Four synthetic event logs, BPIC12, BPIC17,
Pasquadibisceglie et al. (2023) Transductive, heterogeneous Parameter-based, feature-based LSTM model ACT, TS, RES Transfer process knowledge from a prediction/embedding model pre-trained on real-life event data to a new prediction/embedding model fine-tuned on real-life event data, to handle concept drifts and address a prediction task more effectively Embedding model - NAP 12 BPIC event logs
Mehdiyev et al. (2020) Transductive, homogeneous Parameter-based MLP model ACT, RES, DA Transfer process knowledge from an embedding model pre-trained on real event data to a new prediction model fine-tuned on real event data, to address a prediction task more effectively Hash encoding - NAP BPIC12, BPIC13, Helpdesk
Cabrera et al. (2022) Inductive, heterogeneous Feature-based Multi-tasking LSTM model ACT, TS, text Transfer process knowledge from a generic embedding model pre-trained on a large corpus of text data to an embedding model fine-tuned on real-life event data, to address prediction tasks with a successive prediction model more effectively Embedding model Extraction of three time-related features (time passed since previous event, time within the day, time within the week), min-max normalization NAP, NTP BPIC16
Rizk et al. (2023) Transductive, homogeneous Parameter-based, feature-based LSTM model ACT Transfer process knowledge from a prediction model pre-trained on synthetic event data to a prediction model fine-tuned on real-life event data, to address a prediction task more effectively One-hot encoding, embedding layer - NAP Synthetic loan process log, BPIC15, BPIC18
van Luijken et al. (2023) Transductive, homogeneous Parameter-based Multi-tasking GPT/LSTM model with layer freezing ACT, TS Use process knowledge from a prediction model pre-trained on real-life event data to another prediction model fine-tuned on real-life event data, to address a prediction task more effectively One-hot encoding, embedding layer Scalar timestamp, min-max normalization, embedding layer ASP, TSP Helpdesk, RTFM, BPIC12, BPIC13, BPIC15, BPIC17, Sepsis
Nur et al. (2024) Transductive, heterogeneous Parameter-based, feature-based Transformer model with time-related feature block ACT, TS Transfer process knowledge from a prediction model pre-trained on real-life event data to another prediction model fine-tuned on real-life event data, to address a prediction task more effectively Embedding layer, positional encoding Extraction of two time-related features (elapsed time, lagged time) RTP Helpdesk, BPIC20, BPIC12, BPIC11
Ours Transductive, heterogeneous Feature-based, parameter-based LSTM-encoder model with dedicated parameter initialization ACT, TS Transfer process knowledge from a source context to a target context to enable effective PPM between similar processes within an organization or across different organizations Embedding model Extraction of one time-related feature (duration since start), relative scaling OP Helpdesk, BPIC14
Note. ACT = Activity, TS = Timestamp, RES = Resource, DA = Data attribute, OP = Outcome prediction, NAP = Next activity prediction, RTP = Remaining time prediction, ASP = Activity suffix prediction, TSP = Timestamp suffix prediction, BPIC = Business process intelligence challenge.
Furthermore, two PPM techniques combine feature-based and parameter-based TL. Rizk et al. (2023) pre-train an LSTM model on synthetic event data ($\pazocal{D}_{S}$), transfer parameters of this model to a new one, and fine-tune the new model using other synthetic event data ($\pazocal{D}_{T}$) to solve the same prediction task (i.e., next activity prediction). The first layer of the LSTM model is an embedding layer, which enables matching the feature space of $\pazocal{D}_{S}$ to that of $\pazocal{D}_{T}$. The same experiment is performed for the event data of two different real-life event logs. Nur et al. (2024) suggest a technique with a transformer-based DL model for remaining time prediction. They emphasize the use of TL to address the challenge of limited event data. Their approach involved training a base model on a large amount of event data ($\pazocal{D}_{S}$) and fine-tuning it on a smaller amount of event data ($\pazocal{D}_{T}$) to enhance prediction performance for remaining process times. The framework integrates trace encoding, a transferable component block leveraging multi-head self-attention to capture semantic and temporal relationships, and time-related features. Although both techniques incorporate the ideas of feature-based and parameter-based TL in their models, they require a fine-tuning step on event data in the target context. However, this is difficult in intra- or inter-organizational TL settings, where event data are not or very limitedly available in the target context.
Table 1 summarizes the works identified, which use TL for PPM, and shows how they differ from our proposed technique in terms of dimensions, such as transfer goal, different process domains, and fine-tuning in target not necessary.
5 Transfer learning-based technique for predictive process monitoring
This section describes the design of our proposed technique for TL in PPM. The technique is structured into three phases, as shown in Figure 11. In the first phase, initial model building on source, an event log from the source context is loaded, preprocessed, and a DNN-based base model is built and evaluated on preprocessed event data. In the second phase, transfer resources from source to target, relevant resources from the source context (e.g., pre-trained models) are transferred to the target context. If an event log from the target context is available, the base model from the source context can be applied and evaluated in an intermediate step before being used for (ongoing) process instances in the target context. In the third phase, online application of model to target, an ongoing process instance of the target context is received. This instance is then pre-processed using the transferred resources, and the transferred base model is applied to the pre-processed data. Finally, operational support is provided based on the base model’s predictions for process improvement. In the following sections, the three phases of our proposed TL-based technique for PPM are described.
Figure 4: Overview of our TL-based PPM technique. TL-specific and general resources that need to be transferred from the source context $S$ to the target context $T$ are colored blue and gray, respectively.  [^11] The abbreviation “EM” in Figure 11 stands for “embedding model”.
5.1 Initial model building on source
The first phase, which aims to build and evaluate an initial base model $f_{\text{DNN}}(\cdot)$ in the source context $S$, consists of the following six steps.
First, an event log $L_{S}$ is loaded, representing a set of completed traces of a business process from the source context $S$. A complete trace $\sigma$ represents the data for one finished instance of a process, and contains a sequence of events, $e_{1},\dots,e_{N}$, with sequence length $N$. The traces of the event log are assumed to be complete as they enable addressing PPM tasks that refer to the end of a process instance (e.g., process-related outcome prediction). An event $e$ is a tuple $(p,a,t)$, where $p$ is the id of the process instance, $a$ is the process activity, and $t$ is the timestamp. In addition, to retain the temporal order of traces of the event log $L_{S}$, traces are sorted according to the timestamp of their first event.
Second, the activities and the timestamps of the events of the event log $L_{S}$ are encoded. The activities are encoded by transforming their textual values into high-dimensional feature vector representations using a pre-trained embedding model $f_{\text{EMB}}:\sum^{*}\rightarrow\mathrm{R}^{n}$, where $\sum$ is the alphabet, $n$ is the embedding size, and $\sum^{*}$ are all finite activities that can be built from the alphabet $\sum$. For example, common pre-trained models are word2vec (Mikolov et al. 2013), pre-trained global vectors for word representation (GloVe) (Pennington et al. 2014), and BERT (Devlin et al. 2019) models. In doing so, activity names are represented more abstractly than, for example, in the case of a one-hot encoding, facilitating the transfer from a source to a target domain – that is, symmetrical, feature-based TL. Depending on the architecture and capability of the pre-trained embedding model used, the feature vectors representing the values of the activity attribute have different sizes. For example, a word2vec model is a small neural network model with one hidden layer, which learns the relationship between words and produces feature vectors of sizes, such as 100 and 300.121212https://radimrehurek.com/gensim/models/word2vec.html (Accessed 23 July 2025) In contrast, a BERT model is a DNN with several hidden layers (e.g., 12 layers), that learns the meaning of a word in a sentence depending on the surrounding words (context) and produces larger feature vectors, for example, of size 768.131313https://github.com/google-research/bert (Accessed 23 July 2025)
Further, the timestamp of each event is encoded. For that, first, one or more numerical time features are created from the timestamp that are suitable for transfer. An example of such a time feature is duration since start, indicating the temporal difference from the first to the current event of a trace. After feature creation, the values of each feature are divided by a divisor, which is defined for the source domain and has an equivalent meaning in the target domain. For example, such a divisor could be the $x$-th quantile, the mean, or the maximum value of a time feature. With a divisor of this kind, the values of a time feature in the source domain are aligned to the values of the corresponding time feature in the target domain.
Third, prefix encoding is applied to the traces of the event log $L_{S}$. A prefix is a sub-sequence of a trace or the complete trace itself. Prefixes are created to build and apply a base model, which can predict targets (e.g., a process-related outcome) for each time step in a running business process. For the prefix encoding, we apply the index-based sequence encoding approach (Leontjeva et al. 2015). In this approach, all information in the prefix (including the order) is used and features for each attribute of each executed event (index) are generated (Teinemaa et al. 2019). We encode prefixes in this way because we employ an LSTM-based model as base model, which can effectively learn temporal dependencies from the event data of a prefix, from the first to the current event.
Fourth, the prefixes of the event log $L_{S}$ are transformed into the event data $\mathsf{X}_{S}\in\mathrm{R}^{s\times T\times v}$ and the label data $\textbf{y}_{S}\in\mathrm{R}^{s}$ to build a DNN-based base model. Here, $s$ is the number of prefixes, $T$ refers to the number of time steps of the longest prefix, and $v$ denotes the number of features. For the event data $\mathsf{X}_{S}$, the remaining space for prefixes comprising fewer time steps than the prefix with the maximum number of time steps is padded with zeros. For each time step of each prefix, $\mathsf{X}_{S}$ stores information on the activity and the timestamp attribute in the form of extracted features. The concatenation of the activity features and the time features represents the final feature vector with its size corresponding to the number of features $v$ of the event data $\mathsf{X}_{S}$. For the label data, labels are set on the trace level to predict target values as early as possible. This is common in outcome-oriented PPM (Teinemaa et al. 2019), and applicable in our work, as we focus on binary, outcome-based predictions. For example, if the prediction target of interest is the target in-time, the model is applied at each time step of a running process instance to predict whether this instance will be finished in time or not. Finally, the event data $\mathsf{X}_{S}$ and the label data $\mathbf{y}_{S}$ are split into a training set $(\mathsf{X}_{S}^{tr},\mathbf{y}_{S}^{tr})$ and a validation set $(\mathsf{X}_{S}^{val},\mathbf{y}_{S}^{val})$ for model training, and a test set $(\mathsf{X}_{S}^{te},\mathbf{y}_{S}^{te})$ for model evaluation.
Fifth, given the training set of the source domain $(\mathsf{X}_{S}^{tr},\mathbf{y}_{S}^{tr})$, the base model $f_{\text{DNN}}:\mathsf{X}_{i}\in\mathrm{R}^{T\times v}\rightarrow\hat{\mathbf{y}_{i}}\in(0,1)$ is created and trained, mapping event data instances $\mathsf{X}_{i}=(\mathbf{x}_{1},\dots,\mathbf{x}_{T})$, with $x_{t}\in\mathrm{R}^{v}$ onto predictions $\hat{\mathbf{y}}_{i}$. The base model is a two-layer LSTM model. For each layer $l$ = {1,2} and time step $t=1,\dots,T$, the LSTM (Hochreiter and Schmidhuber 1997) transition is
$\mathbf{f}_{g_{t}}^{(l)}$ $=sigmoid\left(\mathbf{U}_{f}^{(l)}\mathbf{h}_{t-1}^{(l)}+\mathbf{W}_{f}^{(l)}\mathbf{x}_{t}^{(l)}+\mathbf{b}_{f}^{(l)}\right)$ (forget gate), (1)
$\mathbf{i}_{g_{t}}^{(l)}$ $=sigmoid\left(\mathbf{U}_{i}^{(l)}\mathbf{h}_{t-1}^{(l)}+\mathbf{W}_{i}^{(l)}\mathbf{x}_{t}^{(l)}+\mathbf{b}_{i}^{(l)}\right)$ (input gate), (2)
$\tilde{\mathbf{c}}_{t}^{(l)}$ $=tanh\left(\mathbf{U}_{g}^{(l)}\mathbf{h}_{(t-1)}^{(l)}+\mathbf{W}_{g}^{(l)}\mathbf{x}_{t}^{(l)}+\mathbf{b}_{g}^{(l)}\right)$ (candidate memory), (3)
$\mathbf{c}_{t}^{(l)}$ $=\mathbf{f}_{g_{t}}^{(l)}\circ\mathbf{c}_{t-1}^{(l)}+\mathbf{i}_{g_{t}}^{(l)}\circ\tilde{\mathbf{c}}_{t}^{(l)}$ (current memory), (4)
$\mathbf{o}_{g_{t}}^{(l)}$ $=sigmoid\left(\mathbf{U}_{o}^{(l)}\mathbf{h}_{t-1}^{(l)}+\mathbf{W}_{o}^{(l)}\mathbf{x}_{t}^{(l)}+\mathbf{b}_{o}^{(l)}\right)$ (output gate), (5)
$\mathbf{h}_{t}^{(l)}$ $=\mathbf{o}_{g_{t}}^{(l)}\circ tanh\left(\mathbf{c}_{t}^{(l)}\right)$ (current hidden state). (6)
In Eqs. (16), $\mathbf{U}_{\bigoplus}$, $\mathbf{W}_{\bigoplus}$, and $\mathbf{b}_{\bigoplus}$ for $\bigoplus\in\{f,i,g,o\}$ are trainable parameters, $\circ$ is the element-wise product, $\mathbf{h}_{t}$ and $\mathbf{c}_{t}$ are the hidden state and cell memory of the LSTM layer $l$. Stacking the two layers yields an LSTM encoder
$(\mathbf{h}_{T}^{(2)},\mathbf{c}_{T}^{(2)})=f_{\text{LSTM}}(\mathsf{X_{i}}),$ (7)
where $\mathbf{h}_{T}^{(2)}$ and $\mathbf{c}_{T}^{(2)}$ is the last hidden state of the second layer. We use two LSTM layers in the base model to learn a multi-level hierarchy of features. The first layer learns features, capturing lower-level, temporal patterns, whereas the second layer, receiving a sequence of these features, can learn features with higher-level abstraction. The learning of such a multi-level hierarchy of features from event data of the source context is necessary in our setting to facilitate the knowledge transfer to the target context.
Then, the hidden state $\mathbf{h}_{T}^{(2)}$ is taken, projected through a fully-connected layer, and a sigmoid activation is applied to the projection’s output to calculate the prediction, as formalized in Eq. (8).
$\mathbf{\hat{y}}=sigmoid(W_{\text{out }}\mathbf{h}_{T}^{(2)}+\mathbf{b}_{\text{out }}).$ (8)
Before base model optimization, the model’s internal parameters are initialized to improve the learning from the event data of the source context and to facilitate knowledge transfer to the target context. All input weight matrices $\mathbf{W}_{\bigoplus}$ for $\bigoplus\in\{f,i,g,o\}$ are Xavier-uniform initialized to center the values around 0, going into sigmoid/tanh activations. This enables the LSTM to start in a flexible area and encourages the optimizer to obtain useful gradients immediately. All recurrent weight matrices $\mathbf{U}_{\bigoplus}$ for $\bigoplus\in\{f,i,g,o\}$ are initialized with a random orthogonal matrix to keep the singular values of the recurrent map near 1 and to reduce the problem of exploiting/vanishing gradients (Arjovsky et al. 2016). In the recurrent part of the LSTM, the recurrent map is the operation that takes the hidden state at t1t-1 and transforms it to the next hidden state at time $t$, using the recurrent weight matrices. Singular values of the recurrent weight matrices measure how much the recurrent map stretches or squashes vectors in different directions. For example, if a singular value is >1>1, that direction grows exponentially over time (i.e., gradients explode). The forget-gate bias $\mathbf{b}_{f}$ is set to +1 (other LSTM biases $\{\mathbf{b}_{i},\mathbf{b}_{g},\mathbf{b}_{o}\}$ are set to 0) to encourage the network to preserve memory at the start of training (Jozefowicz et al. 2015).
For the final layer, the weights $\mathbf{W}_{out}$ are He-uniform initialized with 0 bias to allow for more stability in training, given the fact that our model has one output neuron to model the two possible process-related outcomes (e.g., in-time or not), which our approach assumes.
The internal parameters of the base model for prediction are optimized by solving the following optimization problem:
$\beta^{*}_{S}$ $=\underset{\beta_{S}}{\arg\min}\sum^{s}_{i=1}\pazocal{L}\Big{(}f_{{\text{DNN}}}\Big{(}\mathsf{X}_{S,i}^{tr};\beta_{S}\Big{)},\mathbf{y}_{S,i}^{tr}\Big{)},$ (9)
where $\pazocal{L}$ is a binary cross entropy loss (BCE loss) function, $\beta^{*}_{S}$ are the adjusted weights of the model after training on training event data of the source context $S$, $s$ is the number of prefixes, $\mathsf{X}_{S,i}^{tr}$ and $\mathbf{y}_{S,i}^{tr}$ are the i-th prefix and label of the training event data $\mathsf{X}^{tr}_{S}$ and the training label data $\mathbf{y}_{S}^{tr}$ of the source context $S$. During model optimization, early stopping is performed if no improvement in the BCE loss on the validation event data $(\mathsf{X}_{S}^{val},\mathbf{y}_{S}^{val})$ takes place over 10 subsequent epochs.
Sixth, the built base model $f_{\text{DNN}}(\cdot)$ is applied to the test set $(\mathsf{X}_{S}^{te},\mathbf{y}_{S}^{te})$ and its prediction performance is measured via ML metrics $m\in M$; that is, $m(f_{\text{DNN}}(\mathsf{X}_{S}^{te};\beta^{*}_{S}),\mathbf{y}_{S}^{te}))$. As ML metrics, commonly used metrics in the context of outcome-oriented PPM are calculated, such as the $\text{AUC}_{\text{ROC}}$ and the F1-score (Teinemaa et al. 2019).
5.2 Transfer resources from source to target
The second phase, which aims to transfer resources from the source context $S$ that are relevant for the target context $T$, consists of two steps.
First, relevant resources are transferred from the source context $S$ to the target context $T$. These resources can be grouped into TL-specific resources and general resources. TL-specific resources include the (pre-trained) embedding model for activity encoding, information for timestamp encoding (one or more selected time features for transfer and divisors for feature value scaling), and the pre-trained prediction model for application. General resources include further information on pre-processing (e.g., the type of prefix encoding) and evaluation (e.g., ML metrics for assessing the model’s prediction performance).
Second, if traces in the form of an event log $L_{T}$ are available of the target context $T$, the base model $f_{\text{DNN}}$ from the source context $S$ can be applied and evaluated in an intermediary step before it is used to create predictions based on (ongoing) process instances in the target context $T$. In this intermediary step, first, the event log $L_{T}$ is loaded, and activity and timestamp encoding are performed. For activity encoding, the transferred, pre-trained embedding model $f_{\text{EMB}}(\cdot)$ is applied, whereas for timestamp encoding, the transferred information on time feature selection and scaling is used. After activity and timestamp encoding, prefixes are encoded, the event data $\mathsf{X}_{T}$ and the label data $\mathbf{y}_{T}$ are created, and the data are split into a training set $(\mathsf{X}_{T}^{tr},\mathbf{y}_{T}^{tr})$, a validation set $(\mathsf{X}_{T}^{val},\mathbf{y}_{T}^{val})$, and a test set $(\mathsf{X}_{T}^{te},\mathbf{y}_{T}^{te})$. To ensure that the pre-processing is performed in the same way as in the source context $S$, further pre-processing information is transferred from the source context $S$ to the target context $T$. Moreover, the transferred model $f_{\text{DNN}(\cdot)}$ could be fine-tuned with event data of the target context $T$ (i.e., $(\mathsf{X}_{T}^{tr},\mathbf{y}_{T}^{tr})$ and $(\mathsf{X}_{T}^{val},\mathbf{y}_{T}^{val})$). However, by default, the base model is not fine-tuned with any event data of the target context. Finally, the transferred base model $f_{\text{DNN}}(\cdot)$ is applied to the test set $(\mathsf{X}_{T}^{te},\mathbf{y}_{T}^{te})$ to evaluate its prediction performance using ML metrics $m\in\pazocal{M}$; that is, $m(f_{\text{DNN}}(\mathsf{X}_{T}^{te};\beta^{*}_{S}),\mathbf{y}_{T}^{te})$. For this purpose, evaluation information, such as ML metrics, is transferred from the source to the target context to assess the prediction performance of the base model in the same way as in the source context $S$.
However, for the evaluation in this intermediary step, there are two further aspects to consider. First, as the test sets of the event logs from the target and source contexts are naturally different, the test ML metrics cannot be directly compared with each other. To overcome that, a second prediction model should be trained “from scratch” using the training set $(\mathsf{X}_{T}^{tr},\mathbf{y}_{T}^{tr})$ of the target context $T$. Based on the test set $(\mathsf{X}_{T}^{te},\mathbf{y}_{T}^{te})$, this second prediction model is compared with the transferred base model in terms of prediction performance to evaluate the performance of the transfer. Second, if the event log $L_{T}$ only comprises a few traces, the entire event log should be used as a test set in the evaluation of the base model to get an initial indication of how well the base model performs in the target context.
5.3 Online application of model to target
The third phase, which aims to create predictions for a new, ongoing process instance in the target context $T$ based on the resources transferred from the source context $S$, consists of six steps. First, the (ongoing) process instance $\sigma_{T}$ is loaded, which represents a prefix of a trace that will be completed in the future. Then, activity encoding and timestamp encoding are performed on the (ongoing) process instance. For the activity encoding, the transferred, pre-trained embedding model $f_{\text{EMB}}(\cdot)$ is applied, and for timestamp encoding, the transferred information on time feature selection and scaling is used. In the next step, the prefix is encoded before it is transformed into event data $\mathsf{X}_{T}$. To also ensure that the pre-processing in the online application is performed in the same way as in the source context $S$, further pre-processing information is transferred from the source context $S$ to the target context $T$. Then, the transferred base model $f_{\text{DNN}}(\cdot)$ is applied to the event data $\mathsf{X}_{T}$ to produce the prediction $\hat{y}$. In contrast to the previous two phases, the prediction is not evaluated, as no label information is available in the online application. Instead, in this phase, the prediction is used for operational support by process users or other stakeholders to, for example, proactively draw attention to undesired process outcomes, with the overall objective of improving the performance of the underlying business process.
6 Evaluation
To evaluate our technique, we instantiated it in two real-life use cases. The first one represents an intra-organizational case, whereas the second one is an inter-organizational case. In doing so, we account for the differences between transfer scenarios and can thoroughly evaluate the performance of our technique for both intra- and inter-organizational scenarios. For example, for an intra-organizational use case, to extend the application of an existing prediction model to a new product or service in the same process context, performed activities can be similar, while variations in other dimensions, like sequence of activities or throughput time, can occur. For an inter-organizational use case, the performed activities might even vary in their descriptions.
For each use case, we evaluated the technique using the corresponding instantiation in five regards. First, we evaluated the overall prediction performance of the technique by comparing it to that of traditional PPM techniques. Second, we evaluated the effect of applying different pre-trained models for activity embedding in our technique on the prediction performance. Third, we evaluated the embedding difference between the activities from the source and the target context. Fourth, we evaluated the effect of the proposed relative cross-domain approach for timestamp encoding on the prediction performance compared to other timestamp encoding techniques. Fifth, we evaluated the transfer relevance by comparing the prediction performance of the transferred model with that of models only trained with varying amounts of training event data of the target context. In addition, we conducted further evaluations to assess the prediction performance of our TL-based PPM technique i) when the complete model is used versus sub-variants of it (see Appendix A), ii) when different prefix encoding approaches are applied (see Appendix B), and iii) when activity and timestamp information about prefixes are encoded separately versus jointly (see Appendix C).
6.1 Use case descriptions
6.1.1 Use case I: Intra-organizational transfer
Intra-organizational TL allows an organization to transfer knowledge across its business processes. To evaluate our TL-based PPM technique in the intra-organizational use case, we applied it to process data handling different products. To create this scenario, we used two sub-logs from one organization, specifically sub-logs of the bpic2014 event log [^14] https://data.4tu.nl/articles/_/12692378/1 (Accessed 23 July 2025). The data were provided by an information and communications technologies (ICT) company, Rabobank Group ICT, from the Netherlands. It concerns their ITSM processes, specifically the incident management process. The data were derived from the company’s ITIL service management tool called the HP Service Manager and refer to incidents that are recorded once a reported problem cannot be resolved over the phone. Each incident is related to a service component, that is, one particular product in the bill of materials. The first sub-log WBS72&223, the source log, only includes traces related to the service components WBS000072 and WBS000223. The second sub-log WBS263, our target log, only includes traces related to the service component WBS000263. In the pre-processing, traces longer than 50 events were removed.
Although both event logs are from the same organization, each service component can be handled differently. For example, the distribution of organizational teams assigned to the activities in both logs varies between service components. This different handling is reflected in different event log characteristics in terms of size, number of variants, and activities, which is why the processes are different. The characteristics of both event logs after pre-processing are summarized in Table 2.
Table 2: Characteristics of the event logs used for the intra-organizational use case.
Characteristics
Number of WBS72&223 ($L_{S}$) WBS263 ($L_{T}$)
Instances 2,959 2,204
Instance Variants 1,599 1,476
Events 28,151 26,692
Activities 26 29
Events per Instance* [3; 50; 9.51] [3; 50; 12.11]
Throughput time in days* [0.0002; 167.86; 5.83] [0.0004; 224.94; 4.68]
* [min; max; mean]
An overview of the activities and their distribution can be found in Table 3.
Table 3: Activities and their frequency for both event logs for the intra-organizational use case.
WBS72&223 ($L_{S}$) WBS263 ($L_{T}$)
Activity Frequency Activity Frequency
Assignment 5.336 Assignment 6.376
Reassignment 4.504 Reassignment 5.131
Update 4.259 Update 4.146
Closed 3.343 Operator Update 2.532
Open 2.959 Closed 2.349
Operator Update 2.758 Open 2.204
Status Change 1.953 Caused By CI 2.044
Caused By CI 1.718 Status Change 777
Reopen 320 Description Update 309
Mail to Customer 185 Impact Change 154
Note. Only the ten most frequent activities are listed.
For our evaluation, we predict whether the process instance will be finished in time or not. The label was constructed by taking the 70th quantile of all traces’ duration as the criterion for in-time for each sub-log independently. The value of the criterion for the source log WBS72&223 was 4.94 days, and for the target log WBS263, 4.15 days. Traces with a longer duration were labeled as not in-time.
6.1.2 Use case II: Inter-organizational transfer
In inter-organizational TL, process knowledge is transferred across organizational boundaries. For the evaluation of our technique in the inter-organizational use case, we use two event logs from different organizations. The first one is the bpic2014 event log [^15] https://data.4tu.nl/articles/_/12692378/1 (Accessed 23 July 2025). The second event log called helpdesk [^16] https://data.4tu.nl/articles/dataset/Dataset_belonging_to_the_help_desk_log_of_an_Italian_Company/12675977 (Accessed 23 July 2025) covers the same type of process, an IT ticket management process. Both event logs were derived from a ticketing system and, therefore, are comparable in their structure and notation. In our scenario for TL, the bpic2014 serves as the source log, and the helpdesk event log as the target log. As the event logs from the different organizations exhibit different characteristics in terms of size, number of variants, and activities, the processes are different.
The characteristics of both event logs after the pre-processing are summarized in Table 4. As with the intra-organizational use case, traces longer than 50 events were removed.
Table 4: Characteristics of the event logs used for the inter-organizational use case.
Characteristics
Number of bpic2014 ($L_{S}$) helpdesk ($L_{T}$)
Instances 46,174 4,480
Instance Variants 22,190 226
Events 435,643 21,348
Activities 38 14
Events per Instance* [1; 50; 9.43] [2; 15; 4.66]
Throughput time in days* [0.0; 392.06; 4.87] [30.64; 59.99; 40.86]
* [min; max; mean]
An overview of the activities and their distribution can be found in Table 5.
Table 5: Activities and their frequency for both event logs for the inter-organizational use case.
bpic2014 ($L_{S}$) helpdesk ($L_{T}$)
Activity Frequency Activity Frequency
Assignment 80.123 Take in charge ticket 5.060
Operator Update 50.563 Resolve ticket 4.983
Closed 49.270 Assign seriousness 4.938
Status Change 48.423 Closed 4.574
Open 46.161 Wait 1.463
Reassignment 45.553 Require upgrade 119
Caused By CI 34.090 Insert ticket 118
Update 32.348 Create SW anomaly 67
Quality Indicator Fixed 7.651 Resolve SW anomaly 13
Communication with customer 5.675 Schedule intervention 5
Description Update 4.285 VERIFIED 3
Pending vendor 4.188 RESOLVED 2
External Vendor Assignment 4.145 INVALID 2
Mail to Customer 3.763 DUPLICATE 1
Update from customer 3.235
Note. Only the 15 most frequent activities are listed.
Just like the intra-organizational use case, we predict whether the process instance is in time or not. The label was constructed by taking the 70th quantile of all traces’ duration as the criterion for in-time for each log independently. The value of the criterion for the source log bpic2014 was 2.83 days, and for the target log helpdesk, 44.95 days. Traces with a longer duration were labeled as not in-time.
6.2 Setup
For our first three evaluations on the prediction performance of our technique, we sorted the process instances of the event log of the source context and the target context in temporal order according to the timestamps of their first event. We then split each event log into a 64% training, a 16% validation, and a 20% test set, trained the models, and computed $\text{AUC}_{\text{ROC}}$ and Matthews correlation coefficient (MCC), as well as the weighted precision, recall, and F1-score based on the prediction results on test sets. Among these ML metrics, we choose the $\text{AUC}_{\text{ROC}}$ as the primary metric because it is independent of the threshold, which, therefore, allows for a more holistic evaluation of the models’ prediction performance. We repeated each experiment five times with different seeds and calculated the average values and the standard deviation across the five runs. In addition, for each approach tested in the first three evaluations, we used two additional baselines and compared their prediction performance with the model of the respective approach trained on the event data of the source context and transferred to the target context without fine-tuning. The first baseline is a model trained and tested on the event data of the source context, which serves as a model with a reference prediction performance in the source context. The second baseline is a model trained and tested on the event data of the target context, which serves as a model with a reference prediction performance in the target context.
For the first evaluation, in which we compared the prediction performance of our TL-based technique to that of the traditional PPM approaches, we selected five baseline approaches. The first baseline employs the same LSTM model as in our proposed technique. To encode the activities for this and the other models, one-hot encoding is applied. This is a common technique for encoding categorical attributes of events in PPM (e.g., Rama-Maneiro et al. 2023). However, since we do not know the target context’s activities when training the base model in the source context, only the source context’s activities are used to fit the one-hot encoder. The encoder is applied to transform the target context’s activities. Consequently, if an activity exists in both contexts, it can be encoded in the target context’s event data; otherwise, this is not possible. In addition, the duration since start time feature is min-max scaled using the training event data from the source context. The second baseline employs a logistic regression model with the same encoding strategy as the first approach. The last three baselines use either an extreme gradient boosting (XGBoost), a random forest, or a decision tree model. The activities are one-hot encoded again, while the duration since start time feature is not scaled or normalized. This is because the underlying algorithms for building these models are tree-based and split data based on feature thresholds rather than the absolute magnitude or distribution of features. The models employed in all baselines are commonly used in outcome-oriented PPM (Teinemaa et al. 2019; Maggi et al. 2014; Wang et al. 2019).
For the second evaluation, in which we investigated the effect of using different pre-trained models for activity encoding on prediction performance, three types of embedding models were employed. The first group includes three static word embedding models. The first two models, glove-wiki-gigaword-100 and glove-wiki-gigaword-300, are GloVe models [^17] https://radimrehurek.com/gensim/models/word2vec.html#other-embeddings (Accessed 23 July 2025) that output feature vectors of length 100 and 300 for the activities, respectively. The third model, word2vec-google-news-300, is a pre-trained word2vec model [^18] https://radimrehurek.com/gensim/models/word2vec.html (Accessed 23 July 2025) that encodes activities as vectors of length 300. Although these encoding models have the limitation that all input words must be included in the vocabulary of the pre-trained model, we did not replace any domain-specific acronyms included in the activity names to ensure a fair comparison. For example, we did not replace the acronym “CI” with “configuration item” in the activity name “Caused By CI” of the bpic2014 event log. The second group includes two contextual word embedding models. The first model is bert-base-cased [^19] https://huggingface.co/google-bert/bert-base-cased (Accessed 23 July 2025), a pre-trained BERT model (Devlin et al. 2019). This BERT model has the same architecture as the original BERT model, which is entirely based on the encoder stack of the transformer model (Vaswani et al. 2017). Pre-training was performed on a large corpus using masked language modeling (prediction of randomly masked words in sentences based on the surrounding words) and next sentence prediction objectives. The second model is bert-base-uncased [^20] https://huggingface.co/google-bert/bert-base-uncased (Accessed 23 July 2025), another pre-trained BERT model, which differs from the bert-base-cased model by lower-casing all input text before tokenization. Both models output contextualized word embeddings of length 748 for each token. As activities typically consist of several tokens, a mean pooling operation is performed on top of the contextual word embeddings to produce one embedding vector per activity. The third group includes two contextual sentence embedding models. The first model is all-MiniLM-L12-v2212121https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2 (Accessed 23 July 2025), a fine-tuned version of the pre-trained Mini-LM model (Wang et al. 2020) MiniLM-L12-H384-uncased [^22] https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2 (Accessed 23 July 2025). Fine-tuning is performed using a contrastive learning objective (predicting for a sentence the correct matching sentence from a group of randomly selected sentences), which allows for a better capture of the semantic meaning of entire sentences. The second model is all-mpnet-base-v2232323https://huggingface.co/sentence-transformers/all-mpnet-base-v2 (Accessed 23 July 2025), a version of the pre-trained mpnet model (Song et al. 2020) for which fine-tuning was also performed using a contrastive learning objective. The all-MiniLM-L12-v2 and all-mpnet-base-v2 models output contextualized word embeddings of lengths 384 and 768, respectively, and, as with the contextual word embedding models, a mean pool operation was performed to get one embedding vector per activity.
For the third evaluation, which assesses the quality of the activity embeddings, we examine the similarity of the activity embeddings from the source and target contexts for the intra- and inter-organizational use cases. For this evaluation, we used the embedding models that demonstrated the highest test $\text{AUC}_{\text{ROC}}$ scores in the target context, calculated the Euclidean distance between the embedding vectors, and visualized the distances through a heatmap.
For the fourth evaluation, in which we assessed the effect of the proposed relative cross-domain approach for timestamp encoding on the prediction performance, we employed two different baseline approaches. In the first approach, the time $h$ (in the form of the number of hours), the day of the week $d$, and the month $m$ are extracted from the timestamp and coded as separate features. These variables have a cyclical structure: a Sunday is followed by a Monday, December is followed by January, and the 23rd hour is followed by the 0-th hour. To take this periodic nature into account, the coding is carried out using a sinus function that can depict cyclical relationships. For this purpose, the months are numbered from 0 (January) to 11 (December), the days of the week from 0 (Monday) to 6 (Sunday), and the hours from 0 to 23, and transformed as follows:
$d_{emb}$ $=sin\left(2\cdot\pi\cdot\frac{d}{7}\right),$ (10)
$m_{emb}$ $=sin\left(2\cdot\pi\cdot\frac{m}{12}\right),$ (11)
$h_{emb}$ $=sin\left(2\cdot\pi\cdot\frac{h}{24}\right).$ (12)
This form of coding means, for example, that the values for Sunday and Monday are close together. The advantage of this periodic representation is that cyclical effects, such as seasonal fluctuations or differences between times of day, can be better captured by the model. In the present evaluation, all three features are tested both individually and in combination. The second approach extracts the duration since start time feature from the timestamp, trains an autoencoder model on this feature’s values in the training event data of the source context, and applies the model to all values of this feature in the event data of the source and the target contexts. The idea behind the autoencoder model is to reconstruct the input values using an encoder and a decoder. The encoder is then used to transform the feature values into a compact form (i.e., the latent space).
For the fifth evaluation, we assessed the transfer relevance by comparing the prediction performance of the transferred model with that of models only trained with varying amounts of training event data of the target context. For this evaluation, we used the embedding models that led to the highest test $\text{AUC}_{\text{ROC}}$ scores in the target context.
The LSTM models trained in all experiments have hidden layers with an internal element size of 128. The adaptive moment estimation (Adam) optimizer (Kingma and Ba 2015) was employed for optimizing the models’ internal parameters, with a learning rate of $\text{1e}^{-3}$ and a maximum number of epochs set to 100. All experiments were conducted on a workstation equipped with 12 CPU cores, 128 GB of RAM, and a single NVIDIA RTX 6000 graphics card. All used and produced materials of the experiments can be found in the online repository of this work. [^24] https://github.com/fau-is/tl4ppm
7 Results
7.1 Use case I: Intra-organizational transfer
For the intra-organizational use case, we evaluated the prediction performance of our technique by comparing it with traditional PPM approaches. In addition, we investigated the feature-based transfer of our technique in terms of different activity embeddings, including the quality of embeddings, and different timestamp encodings. Lastly, the transfer relevance was examined.
7.1.1 Prediction performance for transfer learning-based compared to traditional predictive process monitoring
Table 6 shows the prediction performance of the proposed TL-based PPM technique and the two baseline approaches, compared to five traditional PPM approaches, for the intra-organizational use case. We chose the best activity embedding (see Section 7.1.2) and timestamp encoding (see Section 7.1.4) for our proposed technique. It becomes apparent that our proposed technique achieves the highest scores in terms of precision, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$ compared to the traditional PPM approaches employing XGBoost, random forest, decision tree, and LSTM models. The approach using the logistic regression model demonstrates a slightly superior recall and the second-highest F1-score among the evaluated approaches. The prediction performance closest to our proposed TL-based PPM technique in terms of MCC and $\text{AUC}_{\text{ROC}}$ is the traditional PPM technique employing the LSTM model. Comparing the results of the transferred models with the baseline approach Train and test on $L_{T}$, the transferred models of our technique, as well as the techniques using logistic regression, XGBoost, and LSTM models, surpass the baseline approach Train and test on $L_{T}$ in most metrics.
Table 6: Prediction performance of our proposed TL-based PPM technique and traditional PPM techniques for the intra-organizational use case (average and standard deviation over five runs).
PPM technique Precision Recall F1-score MCC $\text{AUC}_{\text{ROC}}$
TL-based PPM Technique
Our proposed technique
Train and test on $L_{S}$ 0.676 ($\pm$.005) 0.702 ($\pm$.022) 0.682 ($\pm$.006) 0.151 ($\pm$.012) 0.641 ($\pm$.012)
Train and test on $L_{T}$ 0.698 ($\pm$.020) 0.707 ($\pm$.020) 0.697 ($\pm$.019) 0.216 ($\pm$.059) 0.688 ($\pm$.021)
Train on $L_{S}$, test on $L_{T}$ 0.707 ($\pm$.010) 0.704 ($\pm$.024) 0.702 ($\pm$.013) 0.241 ($\pm$.027) 0.703 ($\pm$.008)
Traditional PPM Techniques
LSTM
Train and test on $L_{S}$ 0.670 ($\pm$.005) 0.666 ($\pm$.013) 0.668 ($\pm$.008) 0.143 ($\pm$.014) 0.640 ($\pm$.007)
Train and test on $L_{T}$ 0.644 ($\pm$.010) 0.704 ($\pm$.019) 0.655 ($\pm$.013) 0.070 ($\pm$.030) 0.552 ($\pm$.022)
Train on $L_{S}$, test on $L_{T}$ 0.704 ($\pm$.014) 0.654 ($\pm$.016) 0.670 ($\pm$.013) 0.228 ($\pm$.030) 0.683 ($\pm$.022)
Logistic Regression
Train and test on $L_{S}$ 0.681 ($\pm$.000) 0.734 ($\pm$.000) 0.677 ($\pm$.000) 0.136 ($\pm$.000) 0.651 ($\pm$.000)
Train and test on $L_{T}$ 0.648 ($\pm$.000) 0.707 ($\pm$.000) 0.661 ($\pm$.000) 0.081 ($\pm$.000) 0.630 ($\pm$.000)
Train on $L_{S}$, test on $L_{T}$ 0.684 ($\pm$.000) 0.724 ($\pm$.000) 0.690 ($\pm$.000) 0.170 ($\pm$.000) 0.627 ($\pm$.000)
XGBoost
Train and test on $L_{S}$ 0.641 ($\pm$.000) 0.691 ($\pm$.000) 0.657 ($\pm$.000) 0.062 ($\pm$.000) 0.622 ($\pm$.000)
Train and test on $L_{T}$ 0.651 ($\pm$.000) 0.702 ($\pm$.000) 0.664 ($\pm$.000) 0.091 ($\pm$.000) 0.596 ($\pm$.000)
Train on $L_{S}$, test on $L_{T}$ 0.663 ($\pm$.000) 0.701 ($\pm$.000) 0.674 ($\pm$.000) 0.124 ($\pm$.000) 0.599 ($\pm$.000)
Random Forest
Train and test on $L_{S}$ 0.645 ($\pm$.003) 0.672 ($\pm$.003) 0.656 ($\pm$.003) 0.078 ($\pm$.007) 0.599 ($\pm$.004)
Train and test on $L_{T}$ 0.660 ($\pm$.006) 0.696 ($\pm$.005) 0.672 ($\pm$.005) 0.118 ($\pm$.014) 0.617 ($\pm$.004)
Train on $L_{S}$, test on $L_{T}$ 0.653 ($\pm$.007) 0.687 ($\pm$.006) 0.665 ($\pm$.006) 0.101 ($\pm$.018) 0.600 ($\pm$.006)
Decision Tree
Train and test on $L_{S}$ 0.635 ($\pm$.004) 0.621 ($\pm$.006) 0.627 ($\pm$.005) 0.051 ($\pm$.010) 0.548 ($\pm$.005)
Train and test on $L_{T}$ 0.647 ($\pm$.011) 0.632 ($\pm$.013) 0.639 ($\pm$.012) 0.091 ($\pm$.028) 0.546 ($\pm$.016)
Train on $L_{S}$, test on $L_{T}$ 0.641 ($\pm$.002) 0.623 ($\pm$.004) 0.631 ($\pm$.003) 0.075 ($\pm$.006) 0.538 ($\pm$.003)
Note. Best results for training on $L_{S}$ and testing on $L_{T}$ are marked in bold.
7.1.2 Prediction performance for different activity embedding models
To find an effective activity embedding model for our proposed TL-based technique for the intra-organizational use case, we explored seven activity embedding models. Table 7 shows the results of these seven embedding models, which can be divided into three categories: i) static word, ii) contextual word, and iii) contextual sentence embedding models. Embedding the activity feature with the contextual sentence embedding model all-mpnet-base-v2 achieves the best results in all metrics, that is, precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$.
Table 7: Prediction performance of our TL-based PPM technique with the proposed activity embedding created with different pre-trained embedding models for the intra-organizational use case (average and standard deviation over five runs).
Embedding model Precision Recall F1-score MCC $\text{AUC}_{\text{ROC}}$
Static Word Embedding Models
glove-wiki-gigaword-100
Train and test on $L_{S}$ 0.675 ($\pm$.008) 0.676 ($\pm$.051) 0.668 ($\pm$.026) 0.148 ($\pm$.018) 0.634 ($\pm$.017)
Train and test on $L_{T}$ 0.690 ($\pm$.012) 0.699 ($\pm$.021) 0.689 ($\pm$.014) 0.194 ($\pm$.043) 0.671 ($\pm$.024)
Train on $L_{S}$, test on $L_{T}$ 0.697 ($\pm$.021) 0.670 ($\pm$.045) 0.674 ($\pm$.029) 0.209 ($\pm$.047) 0.688 ($\pm$.014)
glove-wiki-gigaword-300
Train and test on $L_{S}$ 0.669 ($\pm$.008) 0.658 ($\pm$.017) 0.663 ($\pm$.012) 0.139 ($\pm$.021) 0.613 ($\pm$.023)
Train and test on $L_{T}$ 0.693 ($\pm$.013) 0.677 ($\pm$.016) 0.684 ($\pm$.014) 0.208 ($\pm$.033) 0.660 ($\pm$.037)
Train on $L_{S}$, test on $L_{T}$ 0.685 ($\pm$.023) 0.642 ($\pm$.027) 0.658 ($\pm$.025) 0.183 ($\pm$.058) 0.656 ($\pm$.036)
word2vec-google-news-300
Train and test on $L_{S}$ 0.661 ($\pm$.013) 0.674 ($\pm$.034) 0.660 ($\pm$.004) 0.113 ($\pm$.033) 0.617 ($\pm$.015)
Train and test on $L_{T}$ 0.706 ($\pm$.022) 0.713 ($\pm$.024) 0.708 ($\pm$.021) 0.240 ($\pm$.058) 0.678 ($\pm$.037)
Train on $L_{S}$, test on $L_{T}$ 0.693 ($\pm$.020) 0.682 ($\pm$.038) 0.678 ($\pm$.016) 0.199 ($\pm$.050) 0.672 ($\pm$.028)
Contextual Word Embedding Models
bert-base-cased
Train and test on $L_{S}$ 0.669 ($\pm$.010) 0.670 ($\pm$.021) 0.667 ($\pm$.008) 0.138 ($\pm$.025) 0.632 ($\pm$.013)
Train and test on $L_{T}$ 0.691 ($\pm$.009) 0.699 ($\pm$.017) 0.693 ($\pm$.011) 0.201 ($\pm$.026) 0.668 ($\pm$.018)
Train on $L_{S}$, test on $L_{T}$ 0.685 ($\pm$.026) 0.661 ($\pm$.041) 0.666 ($\pm$.029) 0.181 ($\pm$.061) 0.667 ($\pm$.043)
bert-base-uncased
Train and test on $L_{S}$ 0.661 ($\pm$.008) 0.680 ($\pm$.016) 0.668 ($\pm$.006) 0.116 ($\pm$.023) 0.607 ($\pm$.014)
Train and test on $L_{T}$ 0.688 ($\pm$.010) 0.685 ($\pm$.034) 0.671 ($\pm$.015) 0.170 ($\pm$.040) 0.644 ($\pm$.016)
Train on $L_{S}$, test on $L_{T}$ 0.674 ($\pm$.028) 0.659 ($\pm$.032) 0.663 ($\pm$.026) 0.156 ($\pm$.072) 0.635 ($\pm$.032)
Contextual Sentence Embedding Models
all-MiniLM-L12-v2
Train and test on $L_{S}$ 0.672 ($\pm$.011) 0.674 ($\pm$.025) 0.672 ($\pm$.017) 0.148 ($\pm$.029) 0.632 ($\pm$.011)
Train and test on $L_{T}$ 0.701 ($\pm$.005) 0.714 ($\pm$.022) 0.703 ($\pm$.009) 0.223 ($\pm$.011) 0.684 ($\pm$.022)
Train on $L_{S}$, test on $L_{T}$ 0.701 ($\pm$.016) 0.657 ($\pm$.018) 0.671 ($\pm$.013) 0.221 ($\pm$.037) 0.693 ($\pm$.015)
all-mpnet-base-v2
Train and test on $L_{S}$ 0.676 ($\pm$.005) 0.702 ($\pm$.022) 0.682 ($\pm$.006) 0.151 ($\pm$.012) 0.641 ($\pm$.012)
Train and test on $L_{T}$ 0.698 ($\pm$.020) 0.707 ($\pm$.020) 0.697 ($\pm$.019) 0.216 ($\pm$.059) 0.688 ($\pm$.021)
Train on $L_{S}$, test on $L_{T}$ 0.707 ($\pm$.010) 0.704 ($\pm$.024) 0.702 ($\pm$.013) 0.241 ($\pm$.027) 0.703 ($\pm$.008)
Note. Best results for training on $L_{S}$ and testing on $L_{T}$ are marked in bold.
7.1.3 Activity embedding quality
To investigate the feature-based TL part of our technique in more detail, we compared the difference in activity embeddings for the all-mpnet-base-v2 embedding model [^25] The figures showing the Euclidean distance between activity embedding vectors created with the other embedding models of our setting can be found in the online repository. (see Figure 5) between $L_{S}$ and $L_{T}$ by calculating the Euclidean distance between the feature vectors.
The source and target logs have 25 activities in common, which can be seen in Figure 5 as the black diagonal line formed by the 0.00 distances between the activity embeddings. In both event logs, $L_{S}$ and $L_{T}$, there are two groups of activities (indicated by the deep-shaded squares in Figure 5), which contain activities that are recorded at a more granular level. The first group deals with communication, either Communication with customer or Communication with vendor. The second group concerns the quality indicator, either referring to Quality Indicator itself, or providing more details, as in Quality Indicator Fixed or Quality Indicator Set. For both groups, the embeddings of the more granular activities are relatively similar, displaying a smaller distance in embedding vectors.
The target log $L_{T}$ has four additional activities that are not represented in the source log $L_{S}$. These are External update, Incident reproduction, OO Response, and Vendor reference. External update has a clear match with Update and Update from customer in $L_{S}$, and Vendor reference with External vendor assignment in $L_{S}$. For Incident reproduction and OO Response, the distances to the activity embeddings of $L_{S}$ are all similarly far apart. No clear match could be found based on the distances in embeddings.
Figure 5: Difference in embedding with model all-mpnet-base-v2 of activities in $L_{S}$ and $L_{T}$ for the intra-organizational use case.
7.1.4 Prediction performance for different timestamp encoding approaches
Next to investigating strategies to embed activities of our TL-based PPM technique, we investigated five additional encoding strategies for the time-based feature duration since start time and compared them to our technique’s relative cross-domain encoding. These five encoding strategies fall into two categories: i) time-based encoding and ii) autoencoder-based encoding. The results in Table 8 indicate a clear trend towards the relative cross-domain encoding as it achieves the highest precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$ out of the six encoding strategies. The second-best encoding in all metrics is the autoencoder-based encoding.
Table 8: Prediction performance of our TL-based PPM technique with the proposed relative cross-domain encoding of the duration since start time feature and other encoding approaches for the intra-organizational use case (average and standard deviation over five runs).
Encoding approach Precision Recall F1-score MCC $\text{AUC}_{\text{ROC}}$
Relative Cross-domain Encoding
Train and test on $L_{S}$ 0.676 ($\pm$.005) 0.702 ($\pm$.022) 0.682 ($\pm$.006) 0.151 ($\pm$.012) 0.641 ($\pm$.012)
Train and test on $L_{T}$ 0.698 ($\pm$.020) 0.707 ($\pm$.020) 0.697 ($\pm$.019) 0.216 ($\pm$.059) 0.688 ($\pm$.021)
Train on $L_{S}$, test on $L_{T}$ 0.707 ($\pm$.010) 0.704 ($\pm$.024) 0.702 ($\pm$.013) 0.241 ($\pm$.027) 0.703 ($\pm$.008)
Time-based Encoding
Hour
Train and test on $L_{S}$ 0.654 ($\pm$.007) 0.672 ($\pm$.016) 0.660 ($\pm$.004) 0.100 ($\pm$.020) 0.612 ($\pm$.013)
Train and test on $L_{T}$ 0.681 ($\pm$.011) 0.685 ($\pm$.031) 0.681 ($\pm$.019) 0.176 ($\pm$.029) 0.642 ($\pm$.032)
Train on $L_{S}$, test on $L_{T}$ 0.685 ($\pm$.012) 0.681 ($\pm$.020) 0.682 ($\pm$.015) 0.188 ($\pm$.031) 0.663 ($\pm$.014)
Weekday
Train and test on $L_{S}$ 0.662 ($\pm$.014) 0.658 ($\pm$.023) 0.658 ($\pm$.013) 0.121 ($\pm$.036) 0.612 ($\pm$.027)
Train and test on $L_{T}$ 0.683 ($\pm$.007) 0.678 ($\pm$.037) 0.675 ($\pm$.018) 0.177 ($\pm$.017) 0.654 ($\pm$.010)
Train on $L_{S}$, test on $L_{T}$ 0.680 ($\pm$.018) 0.665 ($\pm$.031) 0.670 ($\pm$.023) 0.174 ($\pm$.047) 0.657 ($\pm$.033)
Month
Train and test on $L_{S}$ 0.650 ($\pm$.016) 0.661 ($\pm$.048) 0.651 ($\pm$.027) 0.086 ($\pm$.038) 0.592 ($\pm$.022)
Train and test on $L_{T}$ 0.675 ($\pm$.021) 0.693 ($\pm$.015) 0.679 ($\pm$.012) 0.157 ($\pm$.056) 0.630 ($\pm$.041)
Train on $L_{S}$, test on $L_{T}$ 0.679 ($\pm$.019) 0.666 ($\pm$.043) 0.669 ($\pm$.031) 0.170 ($\pm$.047) 0.655 ($\pm$.033)
Hour + Weekday + Month
Train and test on $L_{S}$ 0.667 ($\pm$.009) 0.668 ($\pm$.023) 0.667 ($\pm$.016) 0.134 ($\pm$.023) 0.623 ($\pm$.011)
Train and test on $L_{T}$ 0.655 ($\pm$.017) 0.675 ($\pm$.021) 0.660 ($\pm$.008) 0.106 ($\pm$.044) 0.617 ($\pm$.025)
Train on $L_{S}$, test on $L_{T}$ 0.695 ($\pm$.015) 0.676 ($\pm$.021) 0.682 ($\pm$.014) 0.210 ($\pm$.036) 0.676 ($\pm$.027)
Autoencoder-based Encoding
Train and test on $L_{S}$ 0.665 ($\pm$.013) 0.684 ($\pm$.023) 0.669 ($\pm$.007) 0.125 ($\pm$.036) 0.630 ($\pm$.012)
Train and test on $L_{T}$ 0.692 ($\pm$.014) 0.700 ($\pm$.032) 0.694 ($\pm$.022) 0.204 ($\pm$.036) 0.665 ($\pm$.020)
Train on $L_{S}$, test on $L_{T}$ 0.703 ($\pm$.009) 0.699 ($\pm$.019) 0.698 ($\pm$.009) 0.230 ($\pm$.025) 0.699 ($\pm$.013)
Note. Best results for training on $L_{S}$ and testing on $L_{T}$ are marked in bold.
7.1.5 Transfer relevance
To provide an overview of how much target data should be available to train a model outperforming our transferred model (not fine-tuned), Figure 6 shows the $\text{AUC}_{\text{ROC}}$ for training a model solely on event data available in the target context with the first 1%, 2%, 5%, 10%, 20%, 50%, and 100% of cases using the all-mpnet-base-v2 (see Section 7.1.2) embedding model for activity encoding and the relative cross-domain approach (see Section 7.1.4) for timestamp encoding.
In terms of $\text{AUC}_{\text{ROC}}$, the transferred model achieves a higher $\text{AUC}_{\text{ROC}}$ compared to the majority of the models trained on any amount of available event data of the target context. The prediction performance of the models gradually improves with more training event data. However, our proposed TL-based technique achieves a comparable prediction performance in terms of $\text{AUC}_{\text{ROC}}$ to models trained on the full training event data. These results show that our transferred model can be used in the intra-organizational use case as a considerable substitute for training a separate model in the target context.
Figure 6: Prediction performance of transferred models compared to models trained on increasing training event data in the target context using the pre-trained embedding model all-mpnet-base-v2 for the intra-organizational use case over five runs.
7.2 Use case II: Inter-organizational transfer
For the inter-organizational use case, we proceeded as for the intra-organizational use case and evaluated the prediction performance of our technique compared to traditional PPM approaches, different activity embeddings, as well as embedding quality and time encodings. Furthermore, we examined the transfer relevance by determining the quantity of training event data required in the target context to attain a prediction performance comparable to that of our proposed technique.
7.2.1 Prediction performance for transfer learning-based compared to traditional predictive process monitoring
Table 9 shows the prediction performance of our proposed TL-based PPM technique and the two baseline approaches, compared to five traditional PPM techniques, for the inter-organizational use case. We chose the best activity encoding (see Section 7.2.2) and timestamp encoding (see Section 7.2.4) for our proposed technique. Similar to the intra-organizational use case, our proposed TL-based PPM technique outperforms all traditional PPM approaches in recall, F1 score, and $\text{AUC}_{\text{ROC}}$.
The traditional PPM technique that employs the XGBoost model attains a higher recall, while the approach with the random forest model yields a superior MCC. Comparing the results of the transferred models with the baseline approach Train and test on $L_{T}$, transferred models come close to the prediction performance of the baseline approach Train and test on $L_{T}$ in most metrics.
Table 9: Prediction performance of our proposed TL-based PPM technique and traditional PPM approaches for the inter-organizational use case (average and standard deviation over five runs).
ML approach Precision Recall F1-score MCC $\text{AUC}_{\text{ROC}}$
TL-based PPM Technique
Our proposed technique
Train and test on $L_{S}$ 0.735 ($\pm$.003) 0.694 ($\pm$.003) 0.708 ($\pm$.002) 0.293 ($\pm$.007) 0.730 ($\pm$.005)
Train and test on $L_{T}$ 0.769 ($\pm$.012) 0.769 ($\pm$.037) 0.766 ($\pm$.024) 0.330 ($\pm$.034) 0.775 ($\pm$.018)
Train on $L_{S}$, test on $L_{T}$ 0.739 ($\pm$.020) 0.755 ($\pm$.050) 0.724 ($\pm$.021) 0.206 ($\pm$.029) 0.711 ($\pm$.020)
Traditional PPM Techniques
LSTM
Train and test on $L_{S}$ 0.737 ($\pm$.001) 0.678 ($\pm$.004) 0.695 ($\pm$.004) 0.292 ($\pm$.004) 0.721 ($\pm$.005)
Train and test on $L_{T}$ 0.676 ($\pm$.047) 0.775 ($\pm$.000) 0.678 ($\pm$.001) 0.011 ($\pm$.016) 0.567 ($\pm$.002)
Train on $L_{S}$, test on $L_{T}$ 0.050 ($\pm$.000) 0.224 ($\pm$.000) 0.082 ($\pm$.000) -0.029 ($\pm$.000) 0.435 ($\pm$.012)
Logistic Regression
Train and test on $L_{S}$ 0.708 ($\pm$.000) 0.740 ($\pm$.000) 0.714 ($\pm$.000) 0.218 ($\pm$.000) 0.728 ($\pm$.000)
Train and test on $L_{T}$ 0.759 ($\pm$.000) 0.790 ($\pm$.000) 0.754 ($\pm$.000) 0.269 ($\pm$.000) 0.670 ($\pm$.000)
Train on $L_{S}$, test on $L_{T}$ 0.652 ($\pm$.000) 0.286 ($\pm$.000) 0.231 ($\pm$.000) 0.000 ($\pm$.000) 0.486 ($\pm$.000)
XGBoost
Train and test on $L_{S}$ 0.714 ($\pm$.002) 0.743 ($\pm$.002) 0.720 ($\pm$.001) 0.234 ($\pm$.004) 0.738 ($\pm$.002)
Train and test on $L_{T}$ 0.793 ($\pm$.000) 0.811 ($\pm$.000) 0.793 ($\pm$.000) 0.386 ($\pm$.000) 0.759 ($\pm$.000)
Train on $L_{S}$, test on $L_{T}$ 0.828 ($\pm$.003) 0.257 ($\pm$.044) 0.143 ($\pm$.080) 0.086 ($\pm$.059) 0.604 ($\pm$.004)
Random Forest
Train and test on $L_{S}$ 0.703 ($\pm$.001) 0.710 ($\pm$.001) 0.706 ($\pm$.001) 0.218 ($\pm$.002) 0.692 ($\pm$.001)
Train and test on $L_{T}$ 0.752 ($\pm$.001) 0.756 ($\pm$.002) 0.754 ($\pm$.001) 0.286 ($\pm$.003) 0.724 ($\pm$.002)
Train on $L_{S}$, test on $L_{T}$ 0.827 ($\pm$.002) 0.385 ($\pm$.002) 0.363 ($\pm$.004) 0.227 ($\pm$.001) 0.596 ($\pm$.010)
Decision Tree
Train and test on $L_{S}$ 0.684 ($\pm$.001) 0.681 ($\pm$.001) 0.683 ($\pm$.001) 0.169 ($\pm$.001) 0.590 ($\pm$.001)
Train and test on $L_{T}$ 0.738 ($\pm$.001) 0.726 ($\pm$.001) 0.732 ($\pm$.001) 0.245 ($\pm$.004) 0.676 ($\pm$.003)
Train on $L_{S}$, test on $L_{T}$ 0.794 ($\pm$.002) 0.402 ($\pm$.006) 0.393 ($\pm$.008) 0.204 ($\pm$.006) 0.597 ($\pm$.004)
Note. Best results for training on $L_{S}$ and testing on $L_{T}$ are marked in bold.
7.2.2 Prediction performance for different activity embedding models
In terms of the activity embedding, we again examined seven different embedding models, belonging to the categories static word, contextual word, and contextual embedding models. Table 10 shows the results of these seven embeddings. For the inter-organizational use case, the static word embedding model glove-wiki-gigaword-100 achieves the highest recall, F1-score, and $\text{AUC}_{\text{ROC}}$. The contextual word embedding model bert-base-uncased outperforms other embedding models in terms of precision and MCC.
Table 10: Prediction performance of our TL-based PPM technique with the proposed activity encoding created with different pre-trained embedding models for the inter-organizational use case (average and standard deviation over five runs).
Embedding model Precision Recall F1-score MCC $\text{AUC}_{\text{ROC}}$
Static Word Embedding Models
glove-wiki-gigaword-100
Train and test on $L_{S}$ 0.735 ($\pm$.003) 0.694 ($\pm$.003) 0.708 ($\pm$.002) 0.293 ($\pm$.007) 0.730 ($\pm$.005)
Train and test on $L_{T}$ 0.769 ($\pm$.012) 0.769 ($\pm$.037) 0.766 ($\pm$.024) 0.330 ($\pm$.034) 0.775 ($\pm$.018)
Train on $L_{S}$, test on $L_{T}$ 0.739 ($\pm$.020) 0.755 ($\pm$.050) 0.724 ($\pm$.021) 0.206 ($\pm$.029) 0.711 ($\pm$.020)
glove-wiki-gigaword-300
Train and test on $L_{S}$ 0.738 ($\pm$.003) 0.693 ($\pm$.009) 0.708 ($\pm$.006) 0.300 ($\pm$.005) 0.733 ($\pm$.003)
Train and test on $L_{T}$ 0.756 ($\pm$.025) 0.745 ($\pm$.062) 0.747 ($\pm$.047) 0.294 ($\pm$.074) 0.748 ($\pm$.037)
Train on $L_{S}$, test on $L_{T}$ 0.723 ($\pm$.021) 0.669 ($\pm$.054) 0.685 ($\pm$.040) 0.193 ($\pm$.052) 0.689 ($\pm$.043)
word2vec-google-news-300
Train and test on $L_{S}$ 0.734 ($\pm$.005) 0.701 ($\pm$.010) 0.713 ($\pm$.006) 0.294 ($\pm$.008) 0.732 ($\pm$.004)
Train and test on $L_{T}$ 0.769 ($\pm$.005) 0.774 ($\pm$.014) 0.770 ($\pm$.009) 0.333 ($\pm$.012) 0.784 ($\pm$.005)
Train on $L_{S}$, test on $L_{T}$ 0.735 ($\pm$.030) 0.718 ($\pm$.070) 0.713 ($\pm$.041) 0.215 ($\pm$.073) 0.689 ($\pm$.021)
Contextual Word Embedding Models
bert-base-cased
Train and test on $L_{S}$ 0.738 ($\pm$.002) 0.691 ($\pm$.010) 0.707 ($\pm$.008) 0.300 ($\pm$.003) 0.735 ($\pm$.002)
Train and test on $L_{T}$ 0.676 ($\pm$.042) 0.735 ($\pm$.045) 0.692 ($\pm$.013) 0.084 ($\pm$.050) 0.638 ($\pm$.007)
Train on $L_{S}$, test on $L_{T}$ 0.760 ($\pm$.057) 0.484 ($\pm$.138) 0.483 ($\pm$.142) 0.173 ($\pm$.049) 0.646 ($\pm$.019)
bert-base-uncased
Train and test on $L_{S}$ 0.735 ($\pm$.004) 0.700 ($\pm$.011) 0.713 ($\pm$.008) 0.296 ($\pm$.009) 0.736 ($\pm$.006)
Train and test on $L_{T}$ 0.715 ($\pm$.062) 0.766 ($\pm$.038) 0.727 ($\pm$.052) 0.174 ($\pm$.174) 0.690 ($\pm$.064)
Train on $L_{S}$, test on $L_{T}$ 0.766 ($\pm$.043) 0.618 ($\pm$.173) 0.619 ($\pm$.167) 0.245 ($\pm$.073) 0.692 ($\pm$.042)
Contextual Sentence Embedding Models
all-MiniLM-L12-v2
Train and test on $L_{S}$ 0.736 ($\pm$.003) 0.694 ($\pm$.009) 0.708 ($\pm$.007) 0.296 ($\pm$.006) 0.732 ($\pm$.006)
Train and test on $L_{T}$ 0.772 ($\pm$.007) 0.781 ($\pm$.016) 0.775 ($\pm$.009) 0.342 ($\pm$.016) 0.782 ($\pm$.005)
Train on $L_{S}$, test on $L_{T}$ 0.721 ($\pm$.014) 0.668 ($\pm$.089) 0.681 ($\pm$.069) 0.188 ($\pm$.051) 0.681 ($\pm$.026)
all-mpnet-base-v2
Train and test on $L_{S}$ 0.736 ($\pm$.003) 0.691 ($\pm$.003) 0.706 ($\pm$.002) 0.294 ($\pm$.006) 0.728 ($\pm$.005)
Train and test on $L_{T}$ 0.766 ($\pm$.012) 0.775 ($\pm$.019) 0.769 ($\pm$.014) 0.324 ($\pm$.031) 0.762 ($\pm$.022)
Train on $L_{S}$, test on $L_{T}$ 0.707 ($\pm$.009) 0.550 ($\pm$.040) 0.585 ($\pm$.039) 0.130 ($\pm$.017) 0.655 ($\pm$.014)
Note. Best results for training on $L_{S}$ and testing on $L_{T}$ are marked in bold.
7.2.3 Activity embedding quality
Figure 7 shows the Euclidean distance between the activity embedding vectors from $L_{S}$ and $L_{T}$, created via the glove-wiki-gigaword-100 embedding model. [^26] The figures showing the Euclidean distance between activity embedding vectors created with the other embedding models of our setting can be found in the online repository.
The source and target logs share only two common activities, Closed and RESOLVED, which are represented by an activity embedding distance of 0.00, as shown in Figure 5.
The other twelve activities in $L_{T}$ cannot be found in the source log $L_{S}$. Interestingly, there is no clear match based on the smallest Euclidean distance between the embedding vectors for any of the 12 activities in $L_{T}$ with any of the activities in $L_{S}$. However, based on the embedding distance, these twelve activities can be categorized into two groups. The first group contains seven activities, which have an overall closer match to other activities in $L_{S}$. These are Create SW anomaly, Insert ticket, Require upgrade, Resolve SW anomaly, Resolve ticket, Schedule intervention, and Take in charge ticket. The other five activities in $L_{T}$, Assign seriousness, DUPLICATE, INVALID, VERIFIED, and Wait, are overall further apart from activities in $L_{S}$ based on the Euclidean distance between embedding vectors. However, the activities written in capital letters are outliers in terms of their frequency occurring in $L_{T}$ (see Table 5 in Section 6.1.2).
Figure 7: Difference in embedding with model glove-wiki-gigaword-100 of activities in $L_{S}$ and $L_{T}$ for the inter-organizational use case.
7.2.4 Prediction performance for different timestamp encoding approaches
Similar to the intra-organizational use case, we also examined five additional encoding strategies for the time-based feature duration since start time, belonging to the categories time-based encoding and autoencoder-based encoding, and compared them to our technique’s relative cross-domain encoding. The results in Table 8 indicate a clear trend towards the relative cross-domain encoding as it achieves the highest precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$ out of the six encoding strategies. The prediction performance of our technique using the time-based and autoencoder-based encoding strategies is comparably lower.
Table 11: Prediction performance of our TL-based PPM technique with the proposed cross-domain encoding of the elapsed time attribute and other encoding approaches for the inter-organizational use case (average and standard deviation over five runs).
Encoding approach Precision Recall F1-score MCC $\text{AUC}_{\text{ROC}}$
Relative Cross-domain Encoding
Train and test on $L_{S}$ 0.735 ($\pm$.003) 0.694 ($\pm$.003) 0.708 ($\pm$.002) 0.293 ($\pm$.007) 0.730 ($\pm$.005)
Train and test on $L_{T}$ 0.769 ($\pm$.012) 0.769 ($\pm$.037) 0.766 ($\pm$.024) 0.330 ($\pm$.034) 0.775 ($\pm$.018)
Train on $L_{S}$, test on $L_{T}$ 0.739 ($\pm$.020) 0.755 ($\pm$.050) 0.724 ($\pm$.021) 0.206 ($\pm$.029) 0.711 ($\pm$.020)
Time-based Encoding
Hour
Train and test on $L_{S}$ 0.740 ($\pm$.002) 0.678 ($\pm$.012) 0.696 ($\pm$.010) 0.299 ($\pm$.004) 0.728 ($\pm$.003)
Train and test on $L_{T}$ 0.686 ($\pm$.018) 0.726 ($\pm$.050) 0.687 ($\pm$.007) 0.078 ($\pm$.045) 0.633 ($\pm$.007)
Train on $L_{S}$, test on $L_{T}$ 0.702 ($\pm$.037) 0.657 ($\pm$.126) 0.637 ($\pm$.095) 0.093 ($\pm$.050) 0.609 ($\pm$.013)
Weekday
Train and test on $L_{S}$ 0.739 ($\pm$.004) 0.687 ($\pm$.013) 0.704 ($\pm$.010) 0.300 ($\pm$.009) 0.736 ($\pm$.001)
Train and test on $L_{T}$ 0.686 ($\pm$.013) 0.760 ($\pm$.006) 0.697 ($\pm$.006) 0.072 ($\pm$.026) 0.636 ($\pm$.005)
Train on $L_{S}$, test on $L_{T}$ 0.687 ($\pm$.001) 0.724 ($\pm$.026) 0.698 ($\pm$.004) 0.094 ($\pm$.009) 0.624 ($\pm$.009)
Month
Train and test on $L_{S}$ 0.732 ($\pm$.005) 0.689 ($\pm$.005) 0.703 ($\pm$.003) 0.284 ($\pm$.009) 0.724 ($\pm$.005)
Train and test on $L_{T}$ 0.680 ($\pm$.012) 0.719 ($\pm$.034) 0.690 ($\pm$.005) 0.074 ($\pm$.037) 0.631 ($\pm$.002)
Train on $L_{S}$, test on $L_{T}$ 0.687 ($\pm$.005) 0.674 ($\pm$.047) 0.676 ($\pm$.023) 0.096 ($\pm$.014) 0.628 ($\pm$.005)
Hour + Weekday + Month
Train and test on $L_{S}$ 0.729 ($\pm$.003) 0.707 ($\pm$.016) 0.715 ($\pm$.010) 0.282 ($\pm$.006) 0.724 ($\pm$.005)
Train and test on $L_{T}$ 0.683 ($\pm$.013) 0.737 ($\pm$.031) 0.694 ($\pm$.005) 0.073 ($\pm$.030) 0.628 ($\pm$.010)
Train on $L_{S}$, test on $L_{T}$ 0.681 ($\pm$.002) 0.678 ($\pm$.048) 0.675 ($\pm$.024) 0.080 ($\pm$.007) 0.613 ($\pm$.008)
Autoencoder-based Encoding
Train and test on $L_{S}$ 0.740 ($\pm$.001) 0.684 ($\pm$.007) 0.701 ($\pm$.006) 0.301 ($\pm$.002) 0.732 ($\pm$.004)
Train and test on $L_{T}$ 0.786 ($\pm$.013) 0.801 ($\pm$.016) 0.787 ($\pm$.010) 0.372 ($\pm$.027) 0.785 ($\pm$.006)
Train on $L_{S}$, test on $L_{T}$ 0.732 ($\pm$.058) 0.474 ($\pm$.148) 0.470 ($\pm$.177) 0.135 ($\pm$.082) 0.627 ($\pm$.029)
Note. Best results for training on $L_{S}$ and testing on $L_{T}$ are marked in bold.
7.2.5 Transfer relevance
To provide an overview of how much event data of the target context should be available to train a model outperforming our transferred models (not fine-tuned), Figure 8 shows the $\text{AUC}_{\text{ROC}}$ for training a model solely on event data available in the target context with the first 1%, 2%, 5%, 10%, 20%, 50%, and 100% of cases using the glove-wiki-gigaword-100 (see Section 7.2.2) embedding model for activity encoding and the relative cross-domain approach (see Section 7.2.4) for time encoding.
In terms of $\text{AUC}_{\text{ROC}}$, the transferred model achieves a higher prediction performance compared to the majority of the models trained on 1%, 2%, 5%, 10%, 20% of the available event data of the target context. All models with more than 50% of the event data of the target context outperform our transferred model in terms of $\text{AUC}_{\text{ROC}}$. These results show that our transferred model can be used in the inter-organizational use case as a starting point if event data to train a model in the target context are not readily available in sufficient amounts.
Figure 8: Prediction performance of transferred models compared to models trained on increasing training event data in the target context using the pre-trained embedding model glove-wiki-gigaword-100 for the inter-organizational use case over five runs.
8 Discussion
8.1 Theoretical implications
Our study has multiple implications for research. First, the evaluation results show that the proposed TL-based PPM technique clearly outperforms all the traditional PPM techniques considered in the intra- and inter-organizational use cases in terms of prediction performance. This suggests that a similar understanding of both the source and target contexts is essential for the effective transfer of process knowledge. This is even more important when there is no or very little event data available in the target context for fine-tuning a transferred model. Our work proposes two approaches for aligning the domains of the contexts: i) pre-trained embedding models for activity encoding and ii) a relative cross-domain mapping for timestamp encoding.
Second, our results in the inter- and intra-organizational use cases indicate that pre-trained embedding models can be a better choice to encode the activities for TL in PPM than using a sparse encoding technique, such as one-hot encoding. In addition, depending on the domain (event data) in the source and target contexts, different pre-trained embedding models lead to the best prediction performance of the transferred model in the target context. One possible explanation for this is that there is a correlation between the data used to pre-train the embedding models and the activity attribute values in the event data of the source and target contexts.
Third, we show how embedding models can facilitate TL in PPM. By using these models, process activities from different business processes of the same type, having different names but similar semantic meanings, can be encoded accordingly. This is possible with embedding models because they are trained in a self-supervised manner with auxiliary tasks (e.g., predicting a word in a sentence given the remaining words) to learn the representation of words while capturing their semantics.
Moreover, in our experiments, we observed that two important factors for the transfer of knowledge from one process to another are i) the prediction target and ii) the process information considered for the transfer. Regarding the prediction target, a process-agnostic binary outcome prediction target, like the in-time prediction, is easier for the transfer than a process-specific outcome prediction or the next activity prediction. This is because the latter prediction targets are conceptually more different in the source and the target context. Concerning the considered process information, a transfer that relies solely on the control flow is typically easier than a transfer that relies on the control flow and additional time or data attributes. This is because the consideration of each additional attribute in the event data of the source and target contexts, with its values and distribution, increases the complexity. Nevertheless, to show that information beyond the control flow can be transferred from source to target contexts, we additionally considered the timestamp attribute in the intra- and inter-organization use cases of our evaluation and extracted the time feature duration since start as part of the relative cross-domain encoding.
8.2 Practical implications
Our work also has implications for practice. With a TL-based PPM technique like ours, the adoption of PPM may be fostered in practice. This is specifically beneficial for organizations or areas in organizations without suitable resources, such as computational power, skills or training of employees, or event data for model training. These organizations can use the base model and further resources from another organization that already uses PPM successfully for similar business processes. This allows them to predict, for example, process outcomes or throughput times and, therefore, mitigate risks and decrease costs.
Furthermore, the proposed technique is scalable to incorporate the event data of multiple processes of the same type in the source and target contexts from a conceptual perspective. In this research, we used one event log as a source domain to train our base model, and one event log as a target domain to apply the base model. However, it is conceivable that multiple event logs can be used as a domain in the source context for training our base model. Similarly, multiple event logs can be used as a domain in the target context to apply our base model. To scale our technique in such a way, concepts from the field of federated learning can be incorporated (Verma et al. 2019). As our technique uses an embedding model for activity encoding, scaling is easier than, for example, with a fixed encoding strategy like one-hot encoding, where each activity included in the event data of the source and target contexts has a specific position in the vector.
8.3 Future research
There are four promising directions for future research. First, fine-tuning of the embedding models used in our technique for feature-based TL can be investigated. In doing so, we suspect an improvement in the prediction performance of the model used in the target context. As the source context can comprise multiple business processes and their event data, certain strategies for fine-tuning can be realized to adapt the embedding models to multiple similar processes in the same context in a controlled way. For example, such a strategy could be to randomly take 1,000 traces of each business process’s event log to ensure a fair fine-tuning across multiple business processes with different amounts of event data.
Second, the transfer of information about additional event log dimensions can be investigated. In this paper, we have considered information on the control flow and the time perspectives in the design of our technique. However, information on additional attributes available in the event data, such as data attributes, might be valuable for the transfer in PPM. For this purpose, approaches from the research stream of multidimensional process representation creation and learning in BPM (Weinzierl et al. 2024) can be helpful to find a representation of multiple event log attributes, facilitating the process knowledge transfer from source to target contexts. Closely connected to the information used and its representation lies the question of which information is necessary for the knowledge transfer. This question can be investigated with explainable artificial intelligence (XAI) approaches, which are common in PPM (Stierle et al. 2021).
Third, the application of TL in the context of real business processes of different types can be investigated. In the case of those business processes, the complexity of TL is further increased as the business processes are more different from a structural and semantic perspective. For example, most of the activities, as well as additional data attributes in the event data of a business process in the source context, can differ from the event data of a business process in the target context.
Fourth, the building of process-specific foundation models that are universally usable for various data-driven BPM tasks along the BPM lifecycle (Rizk et al. 2023; Buss et al. 2024) is another promising avenue for future research. For this, existing foundation models for (multivariate) time series can serve as a starting point for model development (e.g., Ye et al. 2024), while available benchmark event logs and other process-relevant resources (e.g., fundamental BPM books, scientific BPM papers, or additional descriptions of event logs) can be used as a data basis for initial model training. As foundational models can support multiple modalities, types of process data going beyond text-transformed event data (e.g., process model collections or process-attached images, audio tracks, or video sequences) are additionally conceivable for model fine-tuning. To fine-tune a foundation model with event data from various organizations, the concept of federated learning can be essential to overcome data privacy restrictions (Verma et al. 2019).
9 Conclusion
In this paper, we proposed a TL-based technique for PPM, allowing organizations without suitable event data or other relevant resources to generate predictions in running business processes for effective decision support. For this purpose, we applied computational DSR (Rai et al. 2017) and designed an artifact in the form of a TL-based technique for PPM, consisting of three phases: i) initial model building on source, ii) transfer base model from source to target, and iii) online application of model to target. By instantiating our technique in the context of intra- and inter-organizational use cases, including event logs for ITSM processes from two organizations, we demonstrated and evaluated its prediction performance. Our results suggest that knowledge of one business process can be transferred to a similar business process in the same or a different organization to enable effective PPM in the target context.
Acknowledgments
Martin Matzner acknowledges funding from the Bavarian State Ministry of Economic Affairs, Regional Development and Energy (StMWi) on “PräMi” (Grant DIK-2307-0002// DIK0533/01) and the German Research Foundation (DFG) on “CoPPA” (Grant 456415646). Furthermore, the authors want to thank Matthias Stierle for his contribution to the initial discussions on the general idea of this research project and Florian Gatzlaff for his support during the project’s early stages.
Appendix A Prediction performance for complete model compared to model sub-variants
We conducted an additional evaluation in the form of an ablation analysis to assess the importance of two main components of our proposed model on prediction performance. These components include the second LSTM layer, which transforms the model into an encoder, and the dedicated parameter initialization, which is performed before model training. When the dedicated parameter initialization is deactivated, all model weights and biases are initialized uniformly. Both model components are designed to improve learning from the event data of the source context and facilitate the transfer of process knowledge to the target context.
Table 12 shows that the complete model of our proposed TL-based technique clearly outperforms its sub-variants in terms of prediction performance when it is trained on $L_{S}$ and tested on $L_{T}$ for the inter-organizational use case. This confirms the importance of using both components in the prediction model for an effective process knowledge transfer from the source to the target contexts.
Table 12: Prediction performance of complete model compared to its sub-variants for the inter-organizational use case (average and standard deviation over five runs).
Scenario Precision Recall F1-score MCC $\text{AUC}_{\text{ROC}}$
Complete Model
Model with two LSTM layers and dedicated parameter initialization
Train and test on $L_{S}$ 0.735 ($\pm$.003) 0.694 ($\pm$.003) 0.708 ($\pm$.002) 0.293 ($\pm$.007) 0.730 ($\pm$.005)
Train and test on $L_{T}$ 0.769 ($\pm$.012) 0.769 ($\pm$.037) 0.766 ($\pm$.024) 0.330 ($\pm$.034) 0.775 ($\pm$.018)
Train on $L_{S}$, test on $L_{T}$ 0.739 ($\pm$.020) 0.755 ($\pm$.050) 0.724 ($\pm$.021) 0.206 ($\pm$.029) 0.711 ($\pm$.020)
Sub-variants of Model
Model with two LSTM layers and without dedicated parameter initialization
Train and test on $L_{S}$ 0.737 ($\pm$.005) 0.691 ($\pm$.009) 0.706 ($\pm$.006) 0.296 ($\pm$.010) 0.730 ($\pm$.005)
Train and test on $L_{T}$ 0.770 ($\pm$.011) 0.783 ($\pm$.015) 0.775 ($\pm$.011) 0.336 ($\pm$.027) 0.767 ($\pm$.015)
Train on $L_{S}$, test on $L_{T}$ 0.712 ($\pm$.037) 0.701 ($\pm$.048) 0.702 ($\pm$.035) 0.167 ($\pm$.099) 0.666 ($\pm$.063)
Model with one LSTM layer and with dedicated parameter initialization
Train and test on $L_{S}$ 0.738 ($\pm$.004) 0.696 ($\pm$.009) 0.710 ($\pm$.007) 0.301 ($\pm$.007) 0.734 ($\pm$.004)
Train and test on $L_{T}$ 0.756 ($\pm$.021) 0.743 ($\pm$.058) 0.745 ($\pm$.043) 0.292 ($\pm$.062) 0.728 ($\pm$.030)
Train on $L_{S}$, test on $L_{T}$ 0.719 ($\pm$.023) 0.694 ($\pm$.046) 0.693 ($\pm$.009) 0.171 ($\pm$.058) 0.701 ($\pm$.022)
Model with one LSTM layer and without dedicated parameter initialization
Train and test on $L_{S}$ 0.739 ($\pm$.005) 0.686 ($\pm$.008) 0.702 ($\pm$.006) 0.299 ($\pm$.009) 0.731 ($\pm$.006)
Train and test on $L_{T}$ 0.764 ($\pm$.016) 0.774 ($\pm$.025) 0.767 ($\pm$.019) 0.318 ($\pm$.045) 0.741 ($\pm$.032)
Train on $L_{S}$, test on $L_{T}$ 0.702 ($\pm$.026) 0.688 ($\pm$.040) 0.687 ($\pm$.017) 0.133 ($\pm$.068) 0.670 ($\pm$.044)
Note. Best results for training on $L_{S}$ and testing on $L_{T}$ are marked in bold.
Appendix B Prediction performance for different prefix encoding approaches
We conducted an additional evaluation to assess the effect of index-based prefix encoding used in our proposed TL-based PPM technique on the prediction performance compared to alternative prefix encoding approaches. The two alternative prefix encoding approaches used in this evaluation were last-3-state encoding and aggregation encoding. Both are common prefix encoding approaches in outcome-oriented PPM (Teinemaa et al. 2019). The first approach, last-3-state encoding, encodes each prefix only by the last three events, or fewer, if the complete trace includes fewer than three events. The second approach, aggregation encoding, encodes each prefix in an aggregated form by calculating the mean value for each attribute across all events in the prefix.
We applied the index-based encoding, the last-3-state encoding, and the aggregation encoding to our proposed TL-based PPM technique for the inter-organizational use case. We also applied these three encoding approaches to two traditional PPM techniques, using LSTM and XGBoost models. We selected these two traditional PPM techniques because the first uses the same LSTM model as in our proposed technique but not in an TL-setting, and the second uses an XGBoost model, which performed relatively well in the first evaluation of our inter-case organizational use case (see Section 7.2.1.
Table 13 shows that our proposed technique with the index-based prefix encoding clearly outperforms all baseline approaches in terms of prediction performance on the test event data of the target context. This suggests that index-based encoding is an effective prefix encoding approach for our proposed TL-based PPM technique.
Table 13: Prediction performance of our proposed TL-based PPM technique and traditional PPM techniques with different prefix encoding approaches for the inter-organizational use case (average and standard deviation over five runs).
ML approach Precision Recall F1-score MCC $\text{AUC}_{\text{ROC}}$
LSTM Model with Activity Embedding and Relative Cross-domain Timestamp Encoding
Index-based encoding
Train and test on $L_{S}$ 0.735 ($\pm$.003) 0.694 ($\pm$.003) 0.708 ($\pm$.002) 0.293 ($\pm$.007) 0.730 ($\pm$.005)
Train and test on $L_{T}$ 0.769 ($\pm$.012) 0.769 ($\pm$.037) 0.766 ($\pm$.024) 0.330 ($\pm$.034) 0.775 ($\pm$.018)
Train on $L_{S}$, test on $L_{T}$ 0.739 ($\pm$.020) 0.755 ($\pm$.050) 0.724 ($\pm$.021) 0.206 ($\pm$.029) 0.711 ($\pm$.020)
Last-3-state encoding
Train and test on $L_{S}$ 0.745 ($\pm$.003) 0.680 ($\pm$.009) 0.698 ($\pm$.007) 0.310 ($\pm$.003) 0.745 ($\pm$.001)
Train and test on $L_{T}$ 0.778 ($\pm$.020) 0.792 ($\pm$.026) 0.779 ($\pm$.017) 0.351 ($\pm$.044) 0.779 ($\pm$.013)
Train on $L_{S}$, test on $L_{T}$ 0.725 ($\pm$.014) 0.694 ($\pm$.054) 0.699 ($\pm$.028) 0.195 ($\pm$.039) 0.709 ($\pm$.025)
Aggregation encoding
Train and test on $L_{S}$ 0.724 ($\pm$.016) 0.697 ($\pm$.015) 0.704 ($\pm$.003) 0.265 ($\pm$.039) 0.708 ($\pm$.025)
Train and test on $L_{T}$ 0.731 ($\pm$.004) 0.741 ($\pm$.018) 0.734 ($\pm$.008) 0.225 ($\pm$.015) 0.706 ($\pm$.013)
Train on $L_{S}$, test on $L_{T}$ 0.718 ($\pm$.023) 0.756 ($\pm$.046) 0.685 ($\pm$.007) 0.074 ($\pm$.015) 0.542 ($\pm$.027)
LSTM Model with One-hot Activity Encoding and Min-max time Feature Encoding
Index-based encoding
Train and test on $L_{S}$ 0.737 ($\pm$.002) 0.677 ($\pm$.004) 0.695 ($\pm$.003) 0.292 ($\pm$.005) 0.721 ($\pm$.005)
Train and test on $L_{T}$ 0.676 ($\pm$.047) 0.775 ($\pm$.000) 0.678 ($\pm$.001) 0.011 ($\pm$.016) 0.567 ($\pm$.002)
Train on $L_{S}$, test on $L_{T}$ 0.050 ($\pm$.000) 0.224 ($\pm$.000) 0.082 ($\pm$.000) -0.029 ($\pm$.000) 0.430 ($\pm$.009)
Last-3-state encoding
Train and test on $L_{S}$ 0.745 ($\pm$.002) 0.675 ($\pm$.005) 0.694 ($\pm$.004) 0.307 ($\pm$.002) 0.741 ($\pm$.002)
Train and test on $L_{T}$ 0.624 ($\pm$.050) 0.776 ($\pm$.000) 0.678 ($\pm$.000) 0.004 ($\pm$.009) 0.556 ($\pm$.004)
Train on $L_{S}$, test on $L_{T}$ 0.143 ($\pm$.208) 0.225 ($\pm$.001) 0.084 ($\pm$.003) -0.007 ($\pm$.015) 0.442 ($\pm$.014)
Aggregation encoding
Train and test on $L_{S}$ 0.725 ($\pm$.002) 0.702 ($\pm$.005) 0.711 ($\pm$.003) 0.271 ($\pm$.004) 0.713 ($\pm$.003)
Train and test on $L_{T}$ 0.602 ($\pm$.000) 0.776 ($\pm$.000) 0.678 ($\pm$.000) 0.000 ($\pm$.000) 0.528 ($\pm$.003)
Train on $L_{S}$, test on $L_{T}$ 0.050 ($\pm$.000) 0.224 ($\pm$.000) 0.082 ($\pm$.000) -0.029 ($\pm$.000) 0.450 ($\pm$.027)
XGBoost with One-hot Activity Encoding
Index-based encoding
Train and test on $L_{S}$ 0.714 ($\pm$.002) 0.743 ($\pm$.002) 0.720 ($\pm$.001) 0.234 ($\pm$.004) 0.738 ($\pm$.002)
Train and test on $L_{T}$ 0.793 ($\pm$.000) 0.811 ($\pm$.000) 0.793 ($\pm$.000) 0.386 ($\pm$.000) 0.759 ($\pm$.000)
Train on $L_{S}$, test on $L_{T}$ 0.828 ($\pm$.003) 0.257 ($\pm$.044) 0.143 ($\pm$.080) 0.086 ($\pm$.059) 0.604 ($\pm$.004)
Last-3-state encoding
Train and test on $L_{S}$ 0.729 ($\pm$.000) 0.750 ($\pm$.000) 0.735 ($\pm$.000) 0.279 ($\pm$.001) 0.759 ($\pm$.000)
Train and test on $L_{T}$ 0.782 ($\pm$.002) 0.801 ($\pm$.002) 0.785 ($\pm$.002) 0.361 ($\pm$.006) 0.763 ($\pm$.002)
Train on $L_{S}$, test on $L_{T}$ 0.707 ($\pm$.003) 0.541 ($\pm$.002) 0.578 ($\pm$.002) 0.128 ($\pm$.005) 0.616 ($\pm$.002)
Aggregation encoding
Train and test on $L_{S}$ 0.723 ($\pm$.000) 0.743 ($\pm$.001) 0.730 ($\pm$.000) 0.265 ($\pm$.001) 0.749 ($\pm$.001)
Train and test on $L_{T}$ 0.736 ($\pm$.002) 0.759 ($\pm$.004) 0.744 ($\pm$.002) 0.236 ($\pm$.004) 0.692 ($\pm$.003)
Train on $L_{S}$, test on $L_{T}$ 0.831 ($\pm$.000) 0.378 ($\pm$.011) 0.351 ($\pm$.016) 0.225 ($\pm$.009) 0.614 ($\pm$.002)
Note. Best results for training on $L_{S}$ and testing on $L_{T}$ are marked in bold.
Appendix C Prediction performance for encoding activity and timestamp information of prefixes separately compared to encoding both jointly
We conducted an additional evaluation to test whether traditional, text-based contextualized embedding models can effectively encode activity and timestamp information about prefixes in a joint manner. For this evaluation, we constructed a baseline technique that jointly encodes the activity and timestamp information about prefixes using one of the four pre-trained contextualized embedding models, which we introduced in the setup of our evaluation. Specifically, the information about a prefix of the length $N$ is fed into the embedding models as $a_{1}$; $t_{1}$; $\dots$; $a_{N}$; $t_{N}$, where $a$ is an activity and $t$ is a timestamp in the format %d-%m-%Y %H:%M:%S.
The encoded prefixes are then mapped via an MLP model onto the values of the in-time prediction target. The MLP model has three fully connected blocks, each of which is followed by tanh activation, batch normalization, and dropout. In addition, the model incorporates residual connections after the second and third blocks using projection layers. The final output is passed through a sigmoid activation function. As with all other DL models used in our evaluation, the internal model parameters were optimized using the Adam optimizer (Kingma and Ba 2015).
Table 14 shows that our TL-based PPM technique achieves a clearly higher prediction performance when trained on $L_{S}$ and tested on $L_{T}$ compared to the baseline techniques using the four different pre-trained contextualized embedding models. This suggests that it is ineffective to use traditional, text-based, contextualized embedding models to jointly encode activity and timestamp information about prefixes in our TL setting.
Table 14: Prediction performance for our TL-based PPM technique that encodes activity and timestamp information about a trace separately compared to a baseline technique that encodes activity and timestamp information about a trace jointly. The baseline technique was tested using the four pre-trained contextualized embeddings of our setup for the inter-organizational use case (average and standard deviation over five runs).
Embedding model Precision Recall F1-score MCC $\text{AUC}_{\text{ROC}}$
Separate Encoding of Activity and Timestamp Information
Proposed approach
Train and test on $L_{S}$ 0.735 ($\pm$.003) 0.694 ($\pm$.003) 0.708 ($\pm$.002) 0.293 ($\pm$.007) 0.730 ($\pm$.005)
Train and test on $L_{T}$ 0.769 ($\pm$.012) 0.769 ($\pm$.037) 0.766 ($\pm$.024) 0.330 ($\pm$.034) 0.775 ($\pm$.018)
Train on $L_{S}$, test on $L_{T}$ 0.739 ($\pm$.020) 0.755 ($\pm$.050) 0.724 ($\pm$.021) 0.206 ($\pm$.029) 0.711 ($\pm$.020)
Joined Encoding of Activity and Timestamp Information
bert-base-cased
Train and test on $L_{S}$ 0.725 ($\pm$.006) 0.591 ($\pm$.031) 0.616 ($\pm$.031) 0.233 ($\pm$.008) 0.680 ($\pm$.005)
Train and test on $L_{T}$ 0.602 ($\pm$.000) 0.776 ($\pm$.000) 0.678 ($\pm$.000) 0.000 ($\pm$.000) 0.517 ($\pm$.020)
Train on $L_{S}$, test on $L_{T}$ 0.688 ($\pm$.006) 0.466 ($\pm$.035) 0.499 ($\pm$.038) 0.075 ($\pm$.016) 0.566 ($\pm$.007)
bert-base-uncased
Train and test on $L_{S}$ 0.714 ($\pm$.008) 0.666 ($\pm$.010) 0.682 ($\pm$.006) 0.238 ($\pm$.016) 0.682 ($\pm$.006)
Train and test on $L_{T}$ 0.602 ($\pm$.000) 0.776 ($\pm$.000) 0.678 ($\pm$.000) 0.000 ($\pm$.000) 0.526 ($\pm$.013)
Train on $L_{S}$, test on $L_{T}$ 0.727 ($\pm$.025) 0.397 ($\pm$.069) 0.393 ($\pm$.101) 0.109 ($\pm$.015) 0.569 ($\pm$.004)
all-MiniLM-L12-v2
Train and test on $L_{S}$ 0.649 ($\pm$.086) 0.668 ($\pm$.078) 0.638 ($\pm$.029) 0.131 ($\pm$.120) 0.598 ($\pm$.096)
Train and test on $L_{T}$ 0.620 ($\pm$.041) 0.750 ($\pm$.057) 0.676 ($\pm$.005) 0.023 ($\pm$.052) 0.537 ($\pm$.055)
Train on $L_{S}$, test on $L_{T}$ 0.645 ($\pm$.040) 0.713 ($\pm$.076) 0.670 ($\pm$.028) 0.036 ($\pm$.035) 0.540 ($\pm$.044)
all-mpnet-base-v2
Train and test on $L_{S}$ 0.651 ($\pm$.087) 0.688 ($\pm$.054) 0.657 ($\pm$.021) 0.140 ($\pm$.131) 0.609 ($\pm$.102)
Train and test on $L_{T}$ 0.656 ($\pm$.031) 0.763 ($\pm$.014) 0.684 ($\pm$.006) 0.021 ($\pm$.021) 0.594 ($\pm$.009)
Train on $L_{S}$, test on $L_{T}$ 0.636 ($\pm$.033) 0.772 ($\pm$.003) 0.679 ($\pm$.002) 0.006 ($\pm$.012) 0.527 ($\pm$.036)
Note. Best results for training on $L_{S}$ and testing on $L_{T}$ are marked in bold.

### v2

@@
 Predictive process monitoring (PPM) transforms these data into value by creating process-related predictions that provide the insights required for proactive interventions at process runtime.
-Existing PPM techniques require sufficient amounts of event data or other relevant resources that might not be readily available, preventing some organizations from utilizing PPM.
+Existing PPM techniques require sufficient amounts of event data or other relevant resources that might not be readily available, which prevents some organizations from utilizing PPM.
 The transfer learning-based PPM technique presented in this paper allows organizations without suitable event data or other relevant resources to implement PPM for effective decision support.
-The technique is instantiated in two real-life use cases, based on which numerical experiments are performed using event logs for IT service management processes in an intra- and inter-organizational setting.
+This technique is instantiated in both a real-life intra- and an inter-organizational use case, based on which numerical experiments are performed using event logs for IT service management processes.
 The results of the experiments suggest that knowledge of one business process can be transferred to a similar business process in the same or a different organization to enable effective PPM in the target context.
-With the proposed technique, organizations can benefit from transfer learning in an intra- and inter-organizational setting, where resources like pre-trained models are transferred within and across organizational boundaries.
+The proposed technique allows organizations to benefit from transfer learning in intra- and inter-organizational settings by transferring resources such as pre-trained models within and across organizational boundaries.
 Keywords: Predictive process monitoring, Process mining, Transfer learning, Language model
@@
 2021).
-Thus, business processes leave large amounts of digital event data reflecting their behavior (van der Aalst 2016).
+Consequently, business processes leave behind large amounts of digital event data that reflect their behavior (van der Aalst 2016).
 To transform these data into value, various business process analytics approaches have been developed (zur Muehlen and Shapiro 2009).
 One of these approaches is predictive process monitoring (PPM) – a set of techniques that leverage historical event log data (Maggi et al.
-2014) to create prediction models Grigori et al.
-(2004), capable of predicting various targets in running business process instances, such as the next activity, remaining time, or process-related outcomes (Di Francescomarino et al.
+2014) to create prediction models (Grigori et al.
+2004), capable of predicting various targets in running business process instances, such as the next activity, remaining time, or process-related outcomes (Di Francescomarino et al.
 2018).
-By generating these process-related predictions, PPM enables proactive interventions in running business process instances, helping to reduce risks and optimize resource allocation (Di Francescomarino et al.
+By creating these process-related predictions, PPM enables proactive interventions in running business process instances, helping to reduce risks and optimize resource allocation (Di Francescomarino et al.
 2018).
@@
 2024).
-These models are trained on historical event log data and then applied to new instances of a running business process in order to predict various targets depending on the business context and analytical goals (Maggi et al.
+These models are trained using historical event log data and then applied to new instances of a running business process in order to predict various targets depending on the business context and the associated analytical goals (Maggi et al.
 2014).
@@
 These include access to sufficient amounts of event log data (Käppel et al.
-2021) and the technical infrastructure required to train complex models, such as graphical processing units.
+2021) and the technical infrastructure required to train complex models, such as graphical processing units (arpteg2018software; Rama-Maneiro et al.
+2023).
 Consequently, some organizations cannot benefit from PPM, while others may already have access to richer event log data or (pre-)trained prediction models for similar business processes.
@@
 The field of transfer learning (TL) provides techniques specifically designed to support such scenarios (Pan and Yang 2009), enabling the transfer of ML models and the knowledge they contain from one learning problem (i.e., domain and task) to another.
-In PPM, TL can be applied in inter-organizational scenarios, in which prediction models are transferred between processes of different organizations, and in intra-organizational scenarios, where a prediction model is transferred from one process to another within a single organization (e.g., across departments or subsidiaries).
+In PPM, TL can be applied in intra-organizational scenarios, in which a prediction model is transferred from one process to another within a single organization (e.g., across departments or subsidiaries), and in inter-organizational scenarios, where prediction models are transferred between processes of different organizations.
 In both scenarios, we consider all resources to be transferred as process knowledge.
 This includes the prediction model (also called the base model).
-Process knowledge originates in its source context, which includes the business process itself, the event data generated by it, and the specific PPM task being addressed using the event data of the business process.
+Process knowledge originates in its source context, which includes the business process itself, the event data produced by it, and the specific PPM task being addressed using the event data of the business process.
 The transferred process knowledge is then applied in the target context (see Figure 1).
-Figure 1: Transfer process knowledge from source to target contexts in PPM.
+Figure 1: Transfer process knowledge from source to target context in PPM.
 Building on this idea, we present a TL-based technique for PPM that relies on existing similarities between business processes and aims to apply a prediction model to a target context.
-Therefore, the prediction model is transferred together with further resources, such as information on the corresponding encoding strategy.
+Therefore, the prediction model is transferred together with further relevant resources, such as information on the corresponding encoding strategy.
 In particular, this work contributes to the existing body of knowledge in four ways:
   1. 1.
-    We present a TL-based technique for PPM that enables the transfer of a base model for an outcome-oriented prediction target without the necessity to fine-tune it on the event data of the target context.
+    We present a TL-based technique for PPM that enables the transfer of a base model for an outcome-oriented prediction target without the necessity to fine-tune it on the event data from the target context.
 Our approach also provides guidance on which resources need to be transferred, including pre-trained models, information on the corresponding encoding strategy, further preprocessing details, and metrics used for measuring the prediction performance.
   2. 2.
-    We propose a cross-domain encoding strategy for event data that enables effective knowledge transfer from source to target contexts.
+    We propose a cross-domain encoding strategy for event data that enables effective knowledge transfer from a source to a target context.
 While timestamp information is encoded via a relative mapping approach, activity information is encoded using pre-trained embedding models to preserve semantic consistency across contexts.
@@
   4. 4.
-    We empirically evaluate our TL-based PPM technique in both an intra- and inter-organizational ITSM use case.
+    We empirically evaluate our TL-based PPM technique in an intra- as well as an inter-organizational ITSM use case.
 The results show that the transferred base model achieves a clearly higher prediction performance than models built with traditional PPM techniques on the event data of the target context.
-To carry out our research, we apply computational design science research (DSR) (Rai et al.
-2017) and structure our work according to the DSR publication schema by Gregor and Hevner (2013): First, we present relevant background on ML and DL in PPM, and TL in general (Section 2).
+To carry out our research, we adopt the computational design science paradigm (Rai et al.
+2017) and structure our work according to the design science research (DSR) publication schema by Gregor and Hevner (2013): First, we present relevant background on ML and DL in PPM, and TL in general (Section 2).
 Second, we explain the practical relevance of TL for PPM in organizations and introduce inter- and intra-organizational TL in the context of PPM (Section 3).
-Subsequently, we give an overview of related work on TL in PPM and position our proposed technique with regard to them (Section 4).
-Then, we provide a detailed description of the proposed artifact (Section 5).
-Afterward, the evaluation of our artifact is described by introducing the intra- and inter-organizational use cases, and providing details on the setup used in the evaluation (Section 6).
-Then, we present the results of our evaluation (Section 7), discuss implications and future research directions (Section 8), and conclude the paper (Section 9).
+Subsequently, we give an overview of related work on TL in PPM and position our proposed technique with regard to it (Section 4).
+We then provide a detailed description of the proposed artifact (Section 5).
+Following this, the evaluation of our artifact is described by introducing the intra- and inter-organizational use case and providing details on the setup used in the evaluation (Section 6).
+Finally, we present the results of our evaluation (Section 7), discuss implications and future research directions (Section 8), and conclude the paper (Section 9).
 2 Background
@@
 130).
-Each event contains at least a case identifier, the name of the executed activity, and a timestamp indicating when the event occurred.
+Every event contains at least a case identifier, the name of the executed activity, and a timestamp indicating when the event occurred.
 Events belonging to the same process instance can be temporally ordered by their timestamp to form a so-called trace (van der Aalst 2016, p.
@@
 2.2 Machine learning in predictive process monitoring
-A branch of early PPM approaches creates prediction models by augmenting discovered process models (or structurally similar representations such as Petri nets or finite automatons) with predictive capabilities (e.g., van der Aalst et al.
-2011; Rogge-Solti and Weske 2015).
-However, as real-world processes are usually more complex than reflected in discovered process models, which frequently fail to capture all execution scenarios, their predictive power is limited (van der Aalst 2011).
-To overcome this limitation, a more recent branch of PPM approaches emerged that exploit the generalization capabilities of ML in identifying and generalizing patterns in event log data to learn prediction models in a data-driven manner.
-Besides a wide range of conventional ML algorithms such as linear regressions, logistic regressions, decision trees, random forests, and gradient boosted decision trees, deep neural networks (DNNs), from the field of DL, have been applied for building prediction models (Marquez-Chamorro et al.
+A branch of early PPM approaches (e.g., van der Aalst et al.
+2011; Rogge-Solti and Weske 2015) create prediction models by augmenting given or discovered process models with predictive capabilities (Oberdorf et al.
+2023).
+These PPM approaches are called process-aware because they use prediction models that exploit an explicit representation of a process model (e.g., a probabilistic finite automaton or a stochastic Petri net) to make a prediction (Marquez-Chamorro et al.
+2017).
+However, as real-world processes are usually more complex than reflected in process models (van der Aalst 2011), the process-model dependence limits the generalizability of the prediction models.
+To overcome this limitation, a more recent branch of PPM approaches (e.g., Maggi et al.
+2014; Tax et al.
+2017) proposes encoding an event log’s traces as feature vectors for the straightforward building of models using ML algorithms (Oberdorf et al.
+2023).
+These approaches are called non-process-aware as the ML models representing the prediction models do not use an explicit description of the process model (Marquez-Chamorro et al.
+2017).
+By leveraging the generalization capabilities of ML models, the approaches often achieve a higher prediction performance than prediction models built on top of process models (Marquez-Chamorro et al.
+2017).
+Besides a wide range of conventional ML algorithms such as linear regressions, logistic regressions, decision trees, random forests, and gradient boosted decision trees, DL algorithms in the form of deep neural networks (DNNs) have been applied for building prediction models (Marquez-Chamorro et al.
 2017; Teinemaa et al.
@@
 2019).
-Especially, the latter has gained prominence due to its ability to learn complex patterns and dependencies in large datasets (LeCun et al.
+They have gained prominence due to their ability to learn complex patterns and dependencies in large datasets (LeCun et al.
 2015).
-DL techniques have shown superior prediction performance and generalizability across various PPM tasks (Kratsch et al.
-2021), and have thus become the state of the art.
-In recent years, a variety of neural network architectures have been employed for different prediction targets, including multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), and recurrent neural network (RNN) with either LSTM cells or gated recurrent units (GRUs), with LSTM networks being favored by the majority (Neu et al.
-2022).
-Among these, RNN, particularly LSTM networks, have been the most widely adopted in PPM.
+Also in PPM, techniques relying on DNNs have shown superior prediction performance and generalizability across various PPM tasks (Kratsch et al.
+2021).
+Therefore, DNNs have become the state-of-the-art in PPM for building prediction models.
+In recent years, a variety of neural network architectures have been employed for building models for different prediction targets, including multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), and recurrent neural networks (RNNs), with either LSTM cells or gated recurrent units (GRUs) (Neu et al.
+2022; Rama-Maneiro et al.
+2023; Weinzierl et al.
+2024).
+Among these, RNNs, particularly LSTM networks, have been the most widely adopted in PPM.
 For example, Evermann et al.
-(2017) applied LSTMs for next activity prediction, while Tax et al.
-(2017) demonstrated that LSTMs outperform conventional ML approaches in predicting both the next activity and its corresponding timestamp.
+(2017) apply LSTM models for the next activity prediction, whereas Tax et al.
+(2017) use LSTM models to predict both the next activity and its corresponding timestamp.
 More recently, transformer-based models, originally introduced in natural language processing, have shown high prediction performance.
@@
 For example, Di Francescomarino et al.
-(2018) proposed multi-perspective models that combine control flow and contextual information for outcome prediction.
+(2018) propose multi-perspective models that combine control-flow and contextual information for outcome prediction.
 As another example, Hinkka et al.
-(2019) suggested clustering categorical event attributes and using the resulting cluster labels as additional information in a RNN.
+(2019) suggest clustering categorical event attributes and using the resulting cluster labels as additional information for a RNN.
 In contrast to these approaches, Pegoraro et al.
@@
 Thus, the encoding determines how transferable and reusable the learned representations are across different but related tasks.
-In many approaches, process instances are transformed into fixed-size feature vectors labeled with a prediction target, as required in supervised ML settings such as classification or regression (Di Francescomarino and Ghidini 2022).
+In many approaches, process instances are transformed into fixed-size feature vectors labeled with a prediction target, as required in supervised ML settings, such as classification or regression (Di Francescomarino and Ghidini 2022).
 However, especially in DL-based models, alternative representations such as event sequences or graph-structured inputs are also employed, allowing models to process richer structural information directly (Weinzierl 2022).
@@
 For the sequence encoding, various strategies have been proposed.
-These strategies range from simple boolean encodings indicating the presence of specific activities to more detailed representations that preserve the full control flow by including the complete sequence of events or the last $n$ events (states) (Di Francescomarino and Ghidini 2022).
+These strategies range from simple boolean encodings, indicating the presence of specific activities, to more detailed representations, preserving the full control flow by including the complete sequence of events or the last $n$ events (states) (Di Francescomarino and Ghidini 2022).
 The encoding of the event attributes depends largely on their data types.
@@
 2019).
-Temporal information, especially timestamp attributes, can be encoded by deriving features, such as time since the last event or the weekday, which then can be treated as numerical or categorical values (Tax et al.
+Temporal information, especially timestamp attributes, can be encoded by deriving features, such as time since the last event or the weekday, which then can be treated as numerical or categorical features (Tax et al.
 2017).
@@
 2016), to extract semantically shallow features from text data.
-By using more advanced techniques that are based on (large) language models, such as contextualized word embedding models like bidirectional encoder representations from transformers (BERT) or robustly optimized BERT pretraining approach (RoBERTa), semantically rich features can be extracted from text data Cabrera et al.
-(2022); Liessmann et al.
-(2024).
+By using more advanced techniques that are based on (large) language models, such as contextualized word embedding models like bidirectional encoder representations from transformers (BERT) or robustly optimized BERT pretraining approach (RoBERTa), semantically rich features can be extracted from text data (Cabrera et al.
+2022; Liessmann et al.
+2024).
 2.3 Transfer learning
@@
 1346).
-A domain $\pazocal{D}=\{\pazocal{X},P(X)\}$ is described by the feature space $\pazocal{X}$, that is, the set of all term vectors, and the marginal probability distribution $P(X)$, where $X$ is an instance set $X=\left\{\mathbf{x}\mid\mathbf{x}_{i}\in\pazocal{X},\text{with }i=1,\dots,n\right\}$.
-Task $\pazocal{T}$ consists of the label space $\pazocal{Y}$, that is, the set of all labels, and the predictive function $f(\cdot)$.
+A domain $\pazocal{D}=\{\pazocal{X},P(X)\}$ is described by the feature space $\pazocal{X}$, that is, the set of all term vectors, and the marginal probability distribution $P(X)$, where $X$ is an instance set $\left\{\mathbf{x}\mid\mathbf{x}_{i}\in\pazocal{X},\text{with }i=1,\dots,n\right\}$.
+Task $\pazocal{T}$ consists of the label space $\pazocal{Y}$, that is, the set of all label values, and the predictive function $f(\cdot)$.
 In TL, a prediction model for the domain with limited data, that is, the target domain, can be improved by transferring knowledge from another domain, the source domain (Weiss et al.
 2016).
-More specifically, the learning of the target prediction function $f_{T}(\cdot)$ for the target learning task $\pazocal{T}_{T}$ and target domain $\pazocal{D}_{T}$ can be improved by transferring knowledge from the source domain $\pazocal{D}_{S}$ and source task $\pazocal{T}_{S}$ with either $\pazocal{D}_{S}\neq\pazocal{D}_{T}$ or $\pazocal{T}_{S}\neq\pazocal{T}_{T}$ (Pan and Yang 2009).
+More specifically, the learning of the target prediction function $f_{T}(\cdot)$, for the target learning task $\pazocal{T}_{T}$, and target domain $\pazocal{D}_{T}$, can be improved by transferring knowledge from the source domain $\pazocal{D}_{S}$, and source task $\pazocal{T}_{S}$, with either $\pazocal{D}_{S}\neq\pazocal{D}_{T}$ or $\pazocal{T}_{S}\neq\pazocal{T}_{T}$ (Pan and Yang 2009).
 In the context of PPM, a domain is represented by the encoded event data $X$.
-The feature space $\pazocal{X}$ of the encoded event data $X$ is determined by the selected event attributes, the type of sequence encoding, and the type of event encoding.
-The marginal distribution $P(X)$ of the encoded event data $X$ is the distribution of the features used to represent the event data $X$.
+The feature space $\pazocal{X}$ of the encoded event data $X$ is determined by the selected event attributes and the used sequence and event coding.
+The marginal probability distribution $P(X)$ of the encoded event data $X$ is the distribution of the features used to represent the event data $X$.
 The addressed PPM task (e.g., the prediction of a process-related outcome) is denoted by $\pazocal{T}$.
-Its goal is to learn a prediction function $f(\cdot)$ that maps a running process instance to a label $y\in\pazocal{Y}$, where $y$ is a value and $\pazocal{Y}$ is the space of all possible values, respectively, for the PPM task $\pazocal{T}$.
+The prediction function $f(\cdot)$ of the task aims to map a running process instance to a label value $y\in\pazocal{Y}$, where $\pazocal{Y}$ is the space of all possible label values for the PPM task $\pazocal{T}$.
 Figure 2: Categories of TL (Zhuang et al. 2020).
@@
 From the problem perspective, depending on the existence of the label, three categories exist: i) transductive TL (label only present in $\pazocal{D}_{S}$, $\pazocal{D}_{S}\neq\pazocal{D}_{T}$, $\pazocal{T}_{S}=\pazocal{T}_{T}$), ii) inductive TL (label present in $\pazocal{D}_{T}$, $\pazocal{T}_{S}\neq\pazocal{T}_{T}$), and iii) unsupervised TL (label unknown in both domains, $\pazocal{T}_{S}\neq\pazocal{T}_{T}$) (Pan and Yang 2009).
-Another differentiation can be made by comparing the feature spaces of the source and target domains.
-While homogeneous TL refers to $\pazocal{X}^{S}=\pazocal{X}^{T}$, in heterogeneous TL, the feature spaces of the source and the target domains are not the same ($\pazocal{X}^{S}\neq\pazocal{X}^{T}$) (Weiss et al.
+Another differentiation can be made by comparing the feature space of the source and the target domain.
+While homogeneous TL refers to $\pazocal{X}^{S}=\pazocal{X}^{T}$, in heterogeneous TL, the feature space of the source and the target domain is not the same ($\pazocal{X}^{S}\neq\pazocal{X}^{T}$) (Weiss et al.
 2016).
@@
 2020).
-In parameter-based approaches, shared (hyper-)parameters or hyperpriors between the source and target domains are transferred.
+In parameter-based approaches, shared (hyper-)parameters or hyperpriors between the source and the target domain are transferred.
 Lastly, relational-based approaches transfer knowledge in the form of relations between the source and target data, assuming that the data are not independent and identically distributed (Pan and Yang 2009).
@@
 Transferring and utilizing knowledge from a source context S to a target context T is particularly interesting in connection with business processes.
-In organizations, IT systems are in place to support the execution of business processes to generate value.
+In organizations, IT systems are in place to support the execution of business processes to create value.
 The widespread use of ERP systems offering similar modules for purchasing, sales and distribution, production, warehouse and inventory management, human resource management, or financial and cost accounting makes business processes comparable across organizations.
 Alongside ERP systems, the use of function-specific software, such as Salesforce for customer relationship management or ServiceNow for IT service management, leads to a standardization of processes.
-In addition to IT systems that offer a common ground for the execution of business processes, organizations can follow industry best practices and standards, for example, by the International Organization for Standardization (ISO)  [^1] https://www.iso.org (Accessed 23 July 2025) or the International Electrotechnical Commission (IEC)  [^2] https://www.iec.ch/homepage (Accessed 23 July 2025).
+In addition to IT systems that offer a common ground for executing business processes, organizations can follow industry best practices and standards.
+For example, they can follow those of the International Organization for Standardization (ISO)  [^1] https://www.iso.org (Accessed 23 July 2025) or the International Electrotechnical Commission (IEC)  [^2] https://www.iec.ch/homepage (Accessed 23 July 2025).
 In IT, standards include ISO/IEC 27001 for information security management systems or the Information Technology Infrastructure Library (ITIL)  [^3] https://www.axelos.com/certifications/itil-service-management/ (Accessed 23 July 2025) for IT service management.
@@
 Besides the voluntary adherence to industry standards, the activities and processes of an organization are also governed by rules and regulations.
-Prime examples of such regulations and directives are General Data Protection Regulation (GDPR)  [^4] https://gdpr.eu (Accessed 23 July 2025) and the recently introduced Corporate Sustainability Reporting Directive (CSRD)  [^5] https://finance.ec.europa.eu (Accessed 23 July 2025).
-Similarly, organizations must follow certain accounting standards based on their location, such as Generally Accepted Accounting Principles (GAAP)  [^6] https://fasab.gov/accounting-standards/ (Accessed 23 July 2025) or International Financial Reporting Standards (IFRS)  [^7] https://www.ifrs.org (Accessed 23 July 2025).
+Prime examples of such regulations and directives are the General Data Protection Regulation (GDPR)  [^4] https://gdpr.eu (Accessed 23 July 2025) and the recently introduced Corporate Sustainability Reporting Directive (CSRD)  [^5] https://finance.ec.europa.eu (Accessed 23 July 2025).
+Similarly, organizations must follow certain accounting standards according to their location, such as the Generally Accepted Accounting Principles (GAAP)  [^6] https://fasab.gov/accounting-standards/ (Accessed 23 July 2025) or the International Financial Reporting Standards (IFRS)  [^7] https://www.ifrs.org (Accessed 23 July 2025).
 Lastly, organizations have similar objectives based on which their processes are designed.
 For example, an organization can strive for cost or time reduction, improved efficiency, increased product or service quality, or customer satisfaction in their process executions.
-The use of comparable software tools, adherence to regulations, and (industry) standards, as well as common objectives, make different business processes within and across organizations increasingly similar to each other.
+The use of comparable software tools, adherence to regulations and (industry) standards, and common objectives make different business processes within and across organizations increasingly similar to each other.
 Figure 3: Intra- and inter-organizational TL for PPM.
@@
 In addition, intra-organizational TL for PPM offers the possibility for an organization to go beyond PPM pilot projects and provide process owners with the tools to implement PPM by themselves.
-Pre-trained prediction models for particular processes can be accessed by process owners to apply PPM without the need for costly resources and extensive model training, almost like an internal model-as-a-service offering.
+Pre-trained prediction models for certain process types can be accessed by process owners to apply PPM without the need for costly resources and extensive model training, almost like an internal model-as-a-service offering.
 While transferring process knowledge through TL within one organization mainly benefits the same organization, TL for PPM can also be applied across organizational boundaries (see Figure 9).
 We coin this inter-organizational TL for PPM.
-For example, a target organization T is interested in the application of PPM and is planning to launch a pilot project to investigate its use for one of its support processes, incident management.
-Organization T is logging all information related to solving incidents in its service desk using ServiceNow, following ITIL while adhering to ISO/IEC 27001.
+For example, a target organization T is interested in the application of PPM and plans to launch a pilot project to investigate its use for one of its support processes, incident management.
+Organization T logs all information related to solving incidents in its service desk using ServiceNow, following ITIL while adhering to ISO/IEC 27001.
 With PPM, organization T aims to reduce the resolution time.
 To keep resources and costs low and to be able to launch a proof-of-concept relatively fast, organization T can use a transferred prediction model, trained with event data of a similar incident management process of source organization S instead of training a prediction model from scratch.
-Another use case for inter-organizational TL for PPM represents the acquisition of organization T by organization S.
+The acquisition of organization T by organization S represents another use case for inter-organizational TL for PPM.
 Some processes of organization S are already supported by PPM.
-As part of the merger, organization S wants to integrate organization T and its ERP system in its PPM efforts.
+As part of the merger, organization S wants to integrate organization T and its ERP system into its PPM efforts.
 To enable smooth integration, the prediction models used for processes in organization S can be transferred to similar processes in organization T.
 The increasing similarities of business processes create an application area for intra- and inter-organizational TL in PPM.
-As inter-organizational TL in PPM aims to transfer process knowledge across organizations, the domains (i.e., the feature space and marginal probability distribution of event data) of the source and target contexts tend to differ more than in intra-organizational TL in PPM, in which knowledge is transferred between business processes within the same organization.
-This greater difference in the source and the target domains in both application cases can be attributed to, for example, organization-specific standards and regulations for performing the same type of business process.
+As inter-organizational TL for PPM aims to transfer process knowledge across organizations, the domains (i.e., the feature space and marginal probability distribution of event data) of the source and target contexts tend to differ more than in intra-organizational TL for PPM, in which knowledge is transferred between business processes within the same organization.
+This greater difference between the source and target domains in an inter-organizational setting can be attributed to, for example, organization-specific standards and regulations for performing the same type of business process.
 4 Related work
-Related TL-based PPM techniques can generally be categorized into two groups: i) PPM techniques that transfer resources within one business process in an organization (i.e., $\pazocal{D}_{S}$ and $\pazocal{D}_{T}$ represent event data on the same business process) and ii) PPM techniques that transfer resources between similar business processes within one organization or across different organizations (i.e., $\pazocal{D}_{S}$ and $\pazocal{D}_{T}$ represent event data on similar business processes within the same or different organizations).
+Related TL-based PPM techniques can generally be categorized into two groups: i) PPM techniques that transfer resources within one business process in an organization (i.e., $\pazocal{D}_{S}$ and $\pazocal{D}_{T}$ represent event data on the same business process), and ii) PPM techniques that transfer resources between similar business processes within one organization or across different organizations (i.e., $\pazocal{D}_{S}$ and $\pazocal{D}_{T}$ represent event data on similar business processes within the same or different organizations).
 Some PPM techniques in the first group aim to transfer parameters between models in transition system-based approaches, to capture patterns in event data better, and ultimately improve prediction performance.
@@
 (2019) suggests a technique for outcome prediction with limited labeled event data.
-An LSTM-encoder model is trained on event data ($\pazocal{D}_{S}$) for next activity prediction, as no extra labels are required for this target.
+An LSTM-encoder model is first trained on event data ($\pazocal{D}_{S}$) for the next activity prediction, as no extra labels are required for this target.
 The last hidden state of the encoder model’s final LSTM layer is then transferred to a second model for outcome prediction.
-This model is then fine-tuned using the same event data ($\pazocal{D}_{T}$) for which a process-related outcome label is available.
+This model is then fine-tuned on a portion of the same event data, for which a process-related outcome label is available ($\pazocal{D}_{T}$).
 Besides parameter-based TL, Cabrera et al.
@@
 This LSTM model addresses the next activity prediction and the next timestamp prediction.
-In addition, the work suggests that fine-tuning the pre-trained BERT models on available event data using the next activity prediction task enhances the LSTM model’s prediction performance for next activities and timestamps as it adapts the representation of the BERT models to the process control-flow.
+In addition, the work suggests that fine-tuning the pre-trained BERT models on available event data using the next activity prediction task enhances the LSTM model’s performance for predicting next activities and timestamps as it adapts the representation of the BERT models to the process control-flow.
 Some PPM techniques in the first group also perform both parameter- and feature-based TL.
@@
 In a similar approach, Chen et al.
-(2022) propose to learn a generic representation of traces by pre-training a BERT on a set of unlabeled trace data ($\pazocal{D}_{S}$) using a custom masked activity modeling task.
-The pre-trained BERT model is then fine-tuned on another set of labeled trace data ($\pazocal{D}_{S}$) for next activity prediction or outcome prediction.
+(2022) propose to learn a generic representation of traces by pre-training a BERT model on a set of unlabeled trace data ($\pazocal{D}_{S}$) using a custom masked activity modeling task.
+The pre-trained BERT model is then fine-tuned on another set of labeled trace data ($\pazocal{D}_{S}$) for the next activity or outcome prediction.
 As in the approach of Pfeiffer et al.
@@
 Pasquadibisceglie et al.
-(2023) suggest using TL for handling concept drift in PPM.
+(2023) suggest using TL for handling concept drifts in PPM.
 Specifically, once a drift is detected within event data ($\pazocal{D}_{S}$), the parameters of both an LSTM model for the next activity prediction and a word2vec model for activity encoding are updated with current event data ($\pazocal{D}_{T}$).
-This improves the prediction performance of the LSTM model and enables the word2vec to deal with newly occurring activities.
-However, even though these PPM techniques incorporate parameter- and feature-based TL, they are not designed to work with different business processes, reflected by general differences in event attribute values and distributions in the source and the target domains, which is necessary for intra- or inter-organizational TL settings.
+This enhances the prediction performance of the LSTM model and enables the word2vec model to deal with newly occurring activities.
+However, even though these PPM techniques incorporate parameter- and feature-based TL, they are not designed to work with different business processes, where the values and distributions of event attributes in the source and target domains are generally different.
+Consequently, these PPM techniques cannot be used in an intra- or inter-organizational TL setting.
 The second group of PPM techniques aims to transfer resources from one business process to another.
@@
 Pfeiffer et al.
-(2021) Inductive, homogeneous Parameter-based, feature-based CNN model with task-specific layer ACT, TS, RES, DA Transfer process knowledge from a model capturing a generic business process representation for various single prediction tasks to prediction models for single tasks to address them more effectively Integer encoding, min-max scaling, Gramian angular field transformation Extraction of one time-related feature (duration since start), min-max normalization, Gramian angular field transformation RTP, NAP, OP Helpdesk, BPIC12, BPIC13, BPIC17, BPIC20, MobIS
+(2021) Inductive, homogeneous Parameter-based, feature-based CNN model with task-specific layer ACT, TS, RES, DA Transfer process knowledge from a model capturing a generic business process representation for various single prediction tasks to prediction models for single tasks, to address them more effectively Integer encoding, min-max scaling, Gramian angular field transformation Extraction of one time-related feature (duration since start), min-max normalization, Gramian angular field transformation RTP, NAP, OP Helpdesk, BPIC12, BPIC13, BPIC17, BPIC20, MobIS
 Chen et al.
-(2022) Inductive, homogeneous Parameter-based, feature-based Multi-tasking, BERT-based transformer model ACT Transfer process knowledge from a generic business process representation for multiple tasks to prediction models for single tasks to address them more effectively Embedding layer, positional encoding - NAP, OP Helpdesk, BPIC12, Sepsis
+(2022) Inductive, homogeneous Parameter-based, feature-based Multi-tasking, BERT-based transformer model ACT Transfer process knowledge from a generic business process representation for multiple tasks to prediction models for single tasks, to address them more effectively Embedding layer, positional encoding - NAP, OP Helpdesk, BPIC12, Sepsis
 Ni et al.
@@
 Cao et al.
-(2022) Transductive, homogeneous Parameter-based MLP model ACT, TS Transfer process knowledge from a prediction model pre-trained on event data (previous transition division) to another prediction model fine-tuned on event data (current transition division), to address a prediction task more effectively One-hot encoding, autoencoder Extraction of eight time-related features (e.g., month, execution time, elapsed time), min-max normalization, autoencoder RTP Four synthetic event logs, BPIC12, BPIC17,
+(2022) Transductive, homogeneous Parameter-based MLP model ACT, TS Transfer process knowledge from a prediction model pre-trained on event data (previous transition division) to another prediction model fine-tuned on event data (current transition division), to address a prediction task more effectively One-hot encoding, autoencoder Extraction of eight time-related features (e.g., month, execution time, and elapsed time), min-max normalization, autoencoder RTP Four synthetic event logs, BPIC12, BPIC17,
 Pasquadibisceglie et al.
@@
 Cabrera et al.
-(2022) Inductive, heterogeneous Feature-based Multi-tasking LSTM model ACT, TS, text Transfer process knowledge from a generic embedding model pre-trained on a large corpus of text data to an embedding model fine-tuned on real-life event data, to address prediction tasks with a successive prediction model more effectively Embedding model Extraction of three time-related features (time passed since previous event, time within the day, time within the week), min-max normalization NAP, NTP BPIC16
+(2022) Inductive, heterogeneous Feature-based Multi-tasking LSTM model ACT, TS, text Transfer process knowledge from a generic embedding model pre-trained on a large corpus of text data to an embedding model fine-tuned on real-life event data, to address prediction tasks with a successive prediction model more effectively Embedding model Extraction of three time-related features (time passed since previous event, time within the day, and time within the week), min-max normalization NAP, NTP BPIC16
 Rizk et al.
@@
 Nur et al.
-(2024) Transductive, heterogeneous Parameter-based, feature-based Transformer model with time-related feature block ACT, TS Transfer process knowledge from a prediction model pre-trained on real-life event data to another prediction model fine-tuned on real-life event data, to address a prediction task more effectively Embedding layer, positional encoding Extraction of two time-related features (elapsed time, lagged time) RTP Helpdesk, BPIC20, BPIC12, BPIC11
+(2024) Transductive, heterogeneous Parameter-based, feature-based Transformer model with time-related feature block ACT, TS Transfer process knowledge from a prediction model pre-trained on real-life event data to another prediction model fine-tuned on real-life event data, to address a prediction task more effectively Embedding layer, positional encoding Extraction of two time-related features (elapsed time and lagged time) RTP Helpdesk, BPIC20, BPIC12, BPIC11
 Ours Transductive, heterogeneous Feature-based, parameter-based LSTM-encoder model with dedicated parameter initialization ACT, TS Transfer process knowledge from a source context to a target context to enable effective PPM between similar processes within an organization or across different organizations Embedding model Extraction of one time-related feature (duration since start), relative scaling OP Helpdesk, BPIC14
@@
 (2024) suggest a technique with a transformer-based DL model for remaining time prediction.
-They emphasize the use of TL to address the challenge of limited event data.
-Their approach involved training a base model on a large amount of event data ($\pazocal{D}_{S}$) and fine-tuning it on a smaller amount of event data ($\pazocal{D}_{T}$) to enhance prediction performance for remaining process times.
-The framework integrates trace encoding, a transferable component block leveraging multi-head self-attention to capture semantic and temporal relationships, and time-related features.
+The transformer-based DL model consists of three components: a trace and positional encoding block, a transferable block (leveraging multi-head self-attention to capture semantic and temporal relationships), and a time-related feature block.
+TL is used in the technique to address the challenge of limited event data.
+This involves training the transformer-based DL model (base model) on a large amount of event data ($\pazocal{D}_{S}$), transferring the trained parameters of its transferable block to a new model, and fine-tuning the new model on a smaller amount of event data ($\pazocal{D}_{T}$) to improve its prediction performance for remaining process times.
 Although both techniques incorporate the ideas of feature-based and parameter-based TL in their models, they require a fine-tuning step on event data in the target context.
 However, this is difficult in intra- or inter-organizational TL settings, where event data are not or very limitedly available in the target context.
-Table 1 summarizes the works identified, which use TL for PPM, and shows how they differ from our proposed technique in terms of dimensions, such as transfer goal, different process domains, and fine-tuning in target not necessary.
+Table 1 summarizes the works identified, which use TL for PPM, and shows how they differ from our proposed technique in terms of dimensions, such as transfer goal, different business processes, and fine-tuning in target not necessary.
 5 Transfer learning-based technique for predictive process monitoring
-This section describes the design of our proposed technique for TL in PPM.
+This section describes the design of our proposed TL-based technique for PPM.
 The technique is structured into three phases, as shown in Figure 11.
@@
 In the third phase, online application of model to target, an ongoing process instance of the target context is received.
-This instance is then pre-processed using the transferred resources, and the transferred base model is applied to the pre-processed data.
+This process instance is then pre-processed using the transferred resources (e.g., information for timestamp encoding), and the transferred base model is applied to the pre-processed data.
 Finally, operational support is provided based on the base model’s predictions for process improvement.
@@
 5.1 Initial model building on source
-The first phase, which aims to build and evaluate an initial base model $f_{\text{DNN}}(\cdot)$ in the source context $S$, consists of the following six steps.
+The first phase, which aims to build and evaluate an initial base model $f_{\text{DNN}}(\cdot)$ in the source context $S$, consists of six steps.
 First, an event log $L_{S}$ is loaded, representing a set of completed traces of a business process from the source context $S$.
@@
 The activities are encoded by transforming their textual values into high-dimensional feature vector representations using a pre-trained embedding model $f_{\text{EMB}}:\sum^{*}\rightarrow\mathrm{R}^{n}$, where $\sum$ is the alphabet, $n$ is the embedding size, and $\sum^{*}$ are all finite activities that can be built from the alphabet $\sum$.
-For example, common pre-trained models are word2vec (Mikolov et al.
-2013), pre-trained global vectors for word representation (GloVe) (Pennington et al.
-2014), and BERT (Devlin et al.
-2019) models.
-In doing so, activity names are represented more abstractly than, for example, in the case of a one-hot encoding, facilitating the transfer from a source to a target domain – that is, symmetrical, feature-based TL.
+For example, word2vec (Mikolov et al.
+2013) and BERT (Devlin et al.
+2019) are common natural-language-processing methods to create pre-trained embedding models.
+By encoding activity names in this way, they are represented more abstractly than, for example, in the case of a one-hot encoding, facilitating the transfer from a source to a target domain – that is, symmetrical, feature-based TL (Zhuang et al.
+2020).
 Depending on the architecture and capability of the pre-trained embedding model used, the feature vectors representing the values of the activity attribute have different sizes.
-For example, a word2vec model is a small neural network model with one hidden layer, which learns the relationship between words and produces feature vectors of sizes, such as 100 and 300.121212https://radimrehurek.com/gensim/models/word2vec.html (Accessed 23 July 2025) In contrast, a BERT model is a DNN with several hidden layers (e.g., 12 layers), that learns the meaning of a word in a sentence depending on the surrounding words (context) and produces larger feature vectors, for example, of size 768.131313https://github.com/google-research/bert (Accessed 23 July 2025)
+For example, a word2vec model is a small neural network model with one hidden layer, which learns the relationship between words and produces feature vectors of lengths, such as 100 and 300.121212https://radimrehurek.com/gensim/models/word2vec.html (Accessed 23 July 2025) In contrast, a BERT model is a DNN with several hidden layers (e.g., 12 layers), that learns the meaning of a word in a sentence depending on the surrounding words (context) and produces larger feature vectors, for example, of length 768.131313https://github.com/google-research/bert (Accessed 23 July 2025)
 Further, the timestamp of each event is encoded.
@@
 This is common in outcome-oriented PPM (Teinemaa et al.
-2019), and applicable in our work, as we focus on binary, outcome-based predictions.
-For example, if the prediction target of interest is the target in-time, the model is applied at each time step of a running process instance to predict whether this instance will be finished in time or not.
+2019), and applicable in our work, as we focus on binary, outcome-based process predictions.
+For example, if the prediction target of interest is in-time, the model is applied at each time step of a running process instance to predict whether this instance will be finished in time or not.
 Finally, the event data $\mathsf{X}_{S}$ and the label data $\mathbf{y}_{S}$ are split into a training set $(\mathsf{X}_{S}^{tr},\mathbf{y}_{S}^{tr})$ and a validation set $(\mathsf{X}_{S}^{val},\mathbf{y}_{S}^{val})$ for model training, and a test set $(\mathsf{X}_{S}^{te},\mathbf{y}_{S}^{te})$ for model evaluation.
-Fifth, given the training set of the source domain $(\mathsf{X}_{S}^{tr},\mathbf{y}_{S}^{tr})$, the base model $f_{\text{DNN}}:\mathsf{X}_{i}\in\mathrm{R}^{T\times v}\rightarrow\hat{\mathbf{y}_{i}}\in(0,1)$ is created and trained, mapping event data instances $\mathsf{X}_{i}=(\mathbf{x}_{1},\dots,\mathbf{x}_{T})$, with $x_{t}\in\mathrm{R}^{v}$ onto predictions $\hat{\mathbf{y}}_{i}$.
+Fifth, given the training event data of the source context $\mathsf{X}_{S}^{tr}=(\mathbf{X}_{1},\dots,\mathbf{X}_{s})$, with $\mathbf{X}_{i}\in\mathrm{R}^{T\times v}$, the base model $f_{\text{DNN}}:\mathbf{X}_{i}\rightarrow\hat{y}_{i}\in(0,1)$ is created and trained, mapping event data instances $\mathbf{X}_{i}=(\mathbf{x}_{1},\dots,\mathbf{x}_{T})$, with $\textbf{x}_{t}\in\mathrm{R}^{v}$, onto estimations $\hat{y}_{i}$.
 The base model is a two-layer LSTM model.
@@
 Stacking the two layers yields an LSTM encoder
-$(\mathbf{h}_{T}^{(2)},\mathbf{c}_{T}^{(2)})=f_{\text{LSTM}}(\mathsf{X_{i}}),$ (7)
+$\Big(\mathbf{h}_{T}^{(2)},\mathbf{c}_{T}^{(2)}\Big)=f_{\text{LSTM}}\Big(\mathbf{X}_{i}\Big),$ (7)
 where $\mathbf{h}_{T}^{(2)}$ and $\mathbf{c}_{T}^{(2)}$ is the last hidden state of the second layer.
@@
 The learning of such a multi-level hierarchy of features from event data of the source context is necessary in our setting to facilitate the knowledge transfer to the target context.
-Then, the hidden state $\mathbf{h}_{T}^{(2)}$ is taken, projected through a fully-connected layer, and a sigmoid activation is applied to the projection’s output to calculate the prediction, as formalized in Eq.
+Then, the hidden state $\mathbf{h}_{T}^{(2)}$ is taken, projected through a fully-connected layer, and a sigmoid activation is applied to the projection’s output to calculate the estimations, as formalized in Eq.
 (8).
-$\mathbf{\hat{y}}=sigmoid(W_{\text{out }}\mathbf{h}_{T}^{(2)}+\mathbf{b}_{\text{out }}).$ (8)
-Before base model optimization, the model’s internal parameters are initialized to improve the learning from the event data of the source context and to facilitate knowledge transfer to the target context.
+$\hat{y}=sigmoid\Big(W_{\text{out }}\mathbf{h}_{T}^{(2)}+\mathbf{b}_{\text{out }}\Big).$ (8)
+Before optimizing the base model’s internal parameters, they are initialized to improve learning from the event data of the source context and to facilitate knowledge transfer to the target context.
 All input weight matrices $\mathbf{W}_{\bigoplus}$ for $\bigoplus\in\{f,i,g,o\}$ are Xavier-uniform initialized to center the values around 0, going into sigmoid/tanh activations.
@@
 2015).
-For the final layer, the weights $\mathbf{W}_{out}$ are He-uniform initialized with 0 bias to allow for more stability in training, given the fact that our model has one output neuron to model the two possible process-related outcomes (e.g., in-time or not), which our approach assumes.
-The internal parameters of the base model for prediction are optimized by solving the following optimization problem:
-$\beta^{*}_{S}$ $=\underset{\beta_{S}}{\arg\min}\sum^{s}_{i=1}\pazocal{L}\Big{(}f_{{\text{DNN}}}\Big{(}\mathsf{X}_{S,i}^{tr};\beta_{S}\Big{)},\mathbf{y}_{S,i}^{tr}\Big{)},$ (9)
-where $\pazocal{L}$ is a binary cross entropy loss (BCE loss) function, $\beta^{*}_{S}$ are the adjusted weights of the model after training on training event data of the source context $S$, $s$ is the number of prefixes, $\mathsf{X}_{S,i}^{tr}$ and $\mathbf{y}_{S,i}^{tr}$ are the i-th prefix and label of the training event data $\mathsf{X}^{tr}_{S}$ and the training label data $\mathbf{y}_{S}^{tr}$ of the source context $S$.
-During model optimization, early stopping is performed if no improvement in the BCE loss on the validation event data $(\mathsf{X}_{S}^{val},\mathbf{y}_{S}^{val})$ takes place over 10 subsequent epochs.
-Sixth, the built base model $f_{\text{DNN}}(\cdot)$ is applied to the test set $(\mathsf{X}_{S}^{te},\mathbf{y}_{S}^{te})$ and its prediction performance is measured via ML metrics $m\in M$; that is, $m(f_{\text{DNN}}(\mathsf{X}_{S}^{te};\beta^{*}_{S}),\mathbf{y}_{S}^{te}))$.
-As ML metrics, commonly used metrics in the context of outcome-oriented PPM are calculated, such as the $\text{AUC}_{\text{ROC}}$ and the F1-score (Teinemaa et al.
+For the final layer, the weights $\mathbf{W}_{out}$ are He-uniform initialized with 0 bias to allow for more stability in training, given the fact that our model has one output neuron to model the two possible process-related outcomes (e.g., in-time or not), which our technique assumes.
+The internal parameters of the base model are optimized by solving the following optimization problem:
+$\beta^{*}_{S}$ $=\underset{\beta_{S}}{\arg\min}\sum^{s}_{i=1}\pazocal{L}\Big(f_{{\text{DNN}}}\Big(\mathbf{X}_{S,i}^{tr};\beta_{S}\Big),y_{S,i}^{tr}\Big),$ (9)
+where $\pazocal{L}$ is a binary cross entropy loss (BCE loss) function, $\beta^{*}_{S}$ are the adjusted internal parameters of the model after training on training event data of the source context $S$, $s$ is the number of prefixes, $\mathbf{X}_{S,i}^{tr}$ is the i-th prefix of the training event data $\mathsf{X}^{tr}_{S}$ of the source context $S$, and $y_{S,i}^{tr}$ is the i-th label value of training label data $\mathbf{y}_{S}^{tr}$ of the source context $S$.
+During the internal parameter optimization, early stopping is performed if no improvement in the BCE loss on the validation set $(\mathsf{X}_{S}^{val},\mathbf{y}_{S}^{val})$ takes place over 10 subsequent epochs.
+Sixth, the built base model $f_{\text{DNN}}(\cdot)$ is evaluated to determine whether it should be transferred from the source context $S$ to the target context $T$.
+To do that, the base model is applied to the test event data $(\mathsf{X}_{S}^{te},\mathbf{y}_{S}^{te})$ and its prediction performance is measured via ML metrics $m\in M$; that is, $m(u(f_{\text{DNN}}(\mathsf{X}_{S}^{te};\beta^{*}_{S})),\mathbf{y}_{S}^{te})$, where $u$ is a function transforming model estimations to class predictions or probabilities.
+As ML metrics, commonly used in outcome-oriented PPM are calculated, such as the $\text{AUC}_{\text{ROC}}$ and the F1-score (Teinemaa et al.
 2019).
+In addition, the prediction performance, measured via the ML metrics, should be considered depending on the used validation strategy, as it determines how the event data is split for model evaluation.
 5.2 Transfer resources from source to target
-The second phase, which aims to transfer resources from the source context $S$ that are relevant for the target context $T$, consists of two steps.
-First, relevant resources are transferred from the source context $S$ to the target context $T$.
+The second phase, which aims to transfer relevant and usable resources from the source context $S$ to the target context $T$, consists of two steps.
+First, relevant resources are transferred from the source context to the target context.
 These resources can be grouped into TL-specific resources and general resources.
-TL-specific resources include the (pre-trained) embedding model for activity encoding, information for timestamp encoding (one or more selected time features for transfer and divisors for feature value scaling), and the pre-trained prediction model for application.
-General resources include further information on pre-processing (e.g., the type of prefix encoding) and evaluation (e.g., ML metrics for assessing the model’s prediction performance).
-Second, if traces in the form of an event log $L_{T}$ are available of the target context $T$, the base model $f_{\text{DNN}}$ from the source context $S$ can be applied and evaluated in an intermediary step before it is used to create predictions based on (ongoing) process instances in the target context $T$.
+TL-specific resources include the pre-trained embedding model for activity encoding, information for timestamp encoding (one or more selected time features for transfer and divisors for feature value scaling), and the pre-trained prediction model for application.
+General resources include further information on pre-processing (e.g., the type of prefix encoding) and evaluation (e.g., ML metrics for measuring the base model’s prediction performance).
+Second, if traces in the form of an event log $L_{T}$ are available of the target context $T$, the base model $f_{\text{DNN}}(\cdot)$ from the source context $S$ can be applied and evaluated in an intermediary step before it is used to create class predictions and probabilities based on (ongoing) process instances in the target context $T$.
 In this intermediary step, first, the event log $L_{T}$ is loaded, and activity and timestamp encoding are performed.
@@
 However, by default, the base model is not fine-tuned with any event data of the target context.
-Finally, the transferred base model $f_{\text{DNN}}(\cdot)$ is applied to the test set $(\mathsf{X}_{T}^{te},\mathbf{y}_{T}^{te})$ to evaluate its prediction performance using ML metrics $m\in\pazocal{M}$; that is, $m(f_{\text{DNN}}(\mathsf{X}_{T}^{te};\beta^{*}_{S}),\mathbf{y}_{T}^{te})$.
-For this purpose, evaluation information, such as ML metrics, is transferred from the source to the target context to assess the prediction performance of the base model in the same way as in the source context $S$.
-However, for the evaluation in this intermediary step, there are two further aspects to consider.
+Finally, the transferred base model is applied to the test set $(\mathsf{X}_{T}^{te},\mathbf{y}_{T}^{te})$ to evaluate its prediction performance.
+For this purpose, evaluation information, such as used ML metrics, is transferred from the source context $S$ to the target context $T$ to assess the prediction performance of the base model in the same way as in the source context.
+For the evaluation in this intermediary step, there are two further aspects to consider.
 First, as the test sets of the event logs from the target and source contexts are naturally different, the test ML metrics cannot be directly compared with each other.
@@
 5.3 Online application of model to target
-The third phase, which aims to create predictions for a new, ongoing process instance in the target context $T$ based on the resources transferred from the source context $S$, consists of six steps.
+The third phase, which aims to create class predictions and probabilities for a new, ongoing process instance in the target context $T$ based on the resources transferred from the source context $S$, consists of six steps.
 First, the (ongoing) process instance $\sigma_{T}$ is loaded, which represents a prefix of a trace that will be completed in the future.
@@
 For the activity encoding, the transferred, pre-trained embedding model $f_{\text{EMB}}(\cdot)$ is applied, and for timestamp encoding, the transferred information on time feature selection and scaling is used.
-In the next step, the prefix is encoded before it is transformed into event data $\mathsf{X}_{T}$.
+In the next step, the prefix is encoded before it is transformed into the event data instance $\mathbf{X}_{T}$.
 To also ensure that the pre-processing in the online application is performed in the same way as in the source context $S$, further pre-processing information is transferred from the source context $S$ to the target context $T$.
-Then, the transferred base model $f_{\text{DNN}}(\cdot)$ is applied to the event data $\mathsf{X}_{T}$ to produce the prediction $\hat{y}$.
-In contrast to the previous two phases, the prediction is not evaluated, as no label information is available in the online application.
-Instead, in this phase, the prediction is used for operational support by process users or other stakeholders to, for example, proactively draw attention to undesired process outcomes, with the overall objective of improving the performance of the underlying business process.
+Then, the transferred base model $f_{\text{DNN}}(\cdot)$ is applied to the event data instance $\mathbf{X}_{T}$ to produce the estimate $\hat{y}$, which is transformed into a class prediction and distribution.
+In contrast to the previous two phases, the prediction performance is not evaluated, as no label information is available in the online application.
+Instead, in this phase, the class prediction and distribution are used for operational support by process users or other stakeholders to, for example, proactively draw attention to undesired process outcomes, with the overall objective of improving the performance of the underlying business process.
 6 Evaluation
 To evaluate our technique, we instantiated it in two real-life use cases.
-The first one represents an intra-organizational case, whereas the second one is an inter-organizational case.
+The first one represents an intra-organizational use case, whereas the second one is an inter-organizational use case.
 In doing so, we account for the differences between transfer scenarios and can thoroughly evaluate the performance of our technique for both intra- and inter-organizational scenarios.
-For example, for an intra-organizational use case, to extend the application of an existing prediction model to a new product or service in the same process context, performed activities can be similar, while variations in other dimensions, like sequence of activities or throughput time, can occur.
-For an inter-organizational use case, the performed activities might even vary in their descriptions.
+For example, in an intra-organizational use case where the application of an existing prediction model is extended to a new product or service within the same process context, the performed activities can be similar, while variations may occur in other dimensions, such as the sequence of activities or throughput time.
+In an inter-organizational use case, the performed activities might even vary in their descriptions.
 For each use case, we evaluated the technique using the corresponding instantiation in five regards.
@@
 To evaluate our TL-based PPM technique in the intra-organizational use case, we applied it to process data handling different products.
-To create this scenario, we used two sub-logs from one organization, specifically sub-logs of the bpic2014 event log [^14] https://data.4tu.nl/articles/_/12692378/1 (Accessed 23 July 2025).
-The data were provided by an information and communications technologies (ICT) company, Rabobank Group ICT, from the Netherlands.
-It concerns their ITSM processes, specifically the incident management process.
-The data were derived from the company’s ITIL service management tool called the HP Service Manager and refer to incidents that are recorded once a reported problem cannot be resolved over the phone.
+In particular, we used the event data of two sub-logs from one organization, specifically sub-logs of the bpic2014 event log [^14] https://data.4tu.nl/articles/_/12692378/1 (Accessed 23 July 2025).
+The data were provided by an information and communications technologies (ICT) company from the Netherlands, called Rabobank Group ICT.
+It concerns their ITSM processes, specifically their incident management process.
+The data were derived from the company’s ITIL service management tool, called the HP Service Manager, and refer to incidents that are recorded when a reported problem cannot be resolved over the phone.
 Each incident is related to a service component, that is, one particular product in the bill of materials.
@@
 An overview of the activities and their distribution can be found in Table 3.
-Table 3: Activities and their frequency for both event logs for the intra-organizational use case.
+Table 3: Activities and their frequency of the event logs used for the intra-organizational use case.
 WBS72&223 ($L_{S}$) WBS263 ($L_{T}$)
@@
 Note. Only the ten most frequent activities are listed.
-For our evaluation, we predict whether the process instance will be finished in time or not.
+For our evaluation, we predict whether running process instances will be finished in time or not.
 The label was constructed by taking the 70th quantile of all traces’ duration as the criterion for in-time for each sub-log independently.
@@
 An overview of the activities and their distribution can be found in Table 5.
-Table 5: Activities and their frequency for both event logs for the inter-organizational use case.
+Table 5: Activities and their frequency of the event logs used for the inter-organizational use case.
 bpic2014 ($L_{S}$) helpdesk ($L_{T}$)
@@
 Note. Only the 15 most frequent activities are listed.
-Just like the intra-organizational use case, we predict whether the process instance is in time or not.
+Just like the intra-organizational use case, we predict whether running process instances will be finished in time or not.
 The label was constructed by taking the 70th quantile of all traces’ duration as the criterion for in-time for each log independently.
@@
 6.2 Setup
-For our first three evaluations on the prediction performance of our technique, we sorted the process instances of the event log of the source context and the target context in temporal order according to the timestamps of their first event.
-We then split each event log into a 64% training, a 16% validation, and a 20% test set, trained the models, and computed $\text{AUC}_{\text{ROC}}$ and Matthews correlation coefficient (MCC), as well as the weighted precision, recall, and F1-score based on the prediction results on test sets.
+For our first three evaluations on the prediction performance of our technique, we sorted the process instances of the event logs of the source and target contexts in temporal order according to the timestamps of their first event.
+We then split each event log into a 64% training, a 16% validation, and a 20% test set, trained the models, and computed the $\text{AUC}_{\text{ROC}}$ and the Matthews correlation coefficient (MCC), as well as the weighted precision, recall, and F1-score based on the prediction results on test sets.
 Among these ML metrics, we choose the $\text{AUC}_{\text{ROC}}$ as the primary metric because it is independent of the threshold, which, therefore, allows for a more holistic evaluation of the models’ prediction performance.
@@
 To encode the activities for this and the other models, one-hot encoding is applied.
-This is a common technique for encoding categorical attributes of events in PPM (e.g., Rama-Maneiro et al.
+It is a common technique for encoding categorical attributes of events in PPM (e.g., Rama-Maneiro et al.
 2023).
 However, since we do not know the target context’s activities when training the base model in the source context, only the source context’s activities are used to fit the one-hot encoder.
-The encoder is applied to transform the target context’s activities.
+The encoder is then applied to transform the target context’s activities.
 Consequently, if an activity exists in both contexts, it can be encoded in the target context’s event data; otherwise, this is not possible.
-In addition, the duration since start time feature is min-max scaled using the training event data from the source context.
+In addition, for the duration since start time feature, a min-max scaler is only fitted on the training event data from the source context and applied to the test event data from the source and target contexts to avoid data leakage.
 The second baseline employs a logistic regression model with the same encoding strategy as the first approach.
@@
 The first group includes three static word embedding models.
-The first two models, glove-wiki-gigaword-100 and glove-wiki-gigaword-300, are GloVe models [^17] https://radimrehurek.com/gensim/models/word2vec.html#other-embeddings (Accessed 23 July 2025) that output feature vectors of length 100 and 300 for the activities, respectively.
+The first two models, glove-wiki-gigaword-100 and glove-wiki-gigaword-300, are pre-trained global vectors for word representation (GloVe) models [^17] https://radimrehurek.com/gensim/models/word2vec.html#other-embeddings (Accessed 23 July 2025) that output feature vectors of length 100 and 300 for the activities, respectively.
 The third model, word2vec-google-news-300, is a pre-trained word2vec model [^18] https://radimrehurek.com/gensim/models/word2vec.html (Accessed 23 July 2025) that encodes activities as vectors of length 300.
@@
 2017).
-Pre-training was performed on a large corpus using masked language modeling (prediction of randomly masked words in sentences based on the surrounding words) and next sentence prediction objectives.
+The pre-training of this model was performed on a large corpus using the masked language modeling (prediction of randomly masked words in sentences based on the surrounding words) and the next sentence prediction objectives.
 The second model is bert-base-uncased [^20] https://huggingface.co/google-bert/bert-base-uncased (Accessed 23 July 2025), another pre-trained BERT model, which differs from the bert-base-cased model by lower-casing all input text before tokenization.
@@
 In the first approach, the time $h$ (in the form of the number of hours), the day of the week $d$, and the month $m$ are extracted from the timestamp and coded as separate features.
-These variables have a cyclical structure: a Sunday is followed by a Monday, December is followed by January, and the 23rd hour is followed by the 0-th hour.
+These variables have a cyclical structure: Sunday is followed by Monday, December is followed by January, and the 23rd hour is followed by the 0-th hour.
 To take this periodic nature into account, the coding is carried out using a sinus function that can depict cyclical relationships.
@@
 For the intra-organizational use case, we evaluated the prediction performance of our technique by comparing it with traditional PPM approaches.
-In addition, we investigated the feature-based transfer of our technique in terms of different activity embeddings, including the quality of embeddings, and different timestamp encodings.
+Furthermore, we investigated the feature-based transfer of our technique in terms of different activity embeddings, including the quality of embeddings, and different timestamp encodings.
 Lastly, the transfer relevance was examined.
@@
 To find an effective activity embedding model for our proposed TL-based technique for the intra-organizational use case, we explored seven activity embedding models.
-Table 7 shows the results of these seven embedding models, which can be divided into three categories: i) static word, ii) contextual word, and iii) contextual sentence embedding models.
-Embedding the activity feature with the contextual sentence embedding model all-mpnet-base-v2 achieves the best results in all metrics, that is, precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$.
+Table 7 shows the results of these models, which can be grouped into three categories: i) static word, ii) contextual word, and iii) contextual sentence embedding models.
+Embedding the activity attribute with the contextual sentence embedding model all-mpnet-base-v2 achieves the best results in all metrics, that is, precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$.
 Table 7: Prediction performance of our TL-based PPM technique with the proposed activity embedding created with different pre-trained embedding models for the intra-organizational use case (average and standard deviation over five runs).
@@
 7.1.3 Activity embedding quality
-To investigate the feature-based TL part of our technique in more detail, we compared the difference in activity embeddings for the all-mpnet-base-v2 embedding model [^25] The figures showing the Euclidean distance between activity embedding vectors created with the other embedding models of our setting can be found in the online repository.
-(see Figure 5) between $L_{S}$ and $L_{T}$ by calculating the Euclidean distance between the feature vectors.
-The source and target logs have 25 activities in common, which can be seen in Figure 5 as the black diagonal line formed by the 0.00 distances between the activity embeddings.
-In both event logs, $L_{S}$ and $L_{T}$, there are two groups of activities (indicated by the deep-shaded squares in Figure 5), which contain activities that are recorded at a more granular level.
+To investigate the feature-based TL part of our technique in more detail, we compared the difference in activity embeddings for the all-mpnet-base-v2 embedding model [^25] The figures showing the Euclidean distance between activity embedding vectors created with the other embedding models of our setting can be found in the online repository of this work.
+between $L_{S}$ and $L_{T}$ by calculating the Euclidean distance between the embedding vectors (see Figure 5).
+The source and target logs have 25 activities in common, which can be seen by the black diagonal line formed by the 0.00 distances between the activity embeddings.
+In addition, in both event logs, $L_{S}$ and $L_{T}$, there are two groups of activities (indicated by the deep-shaded squares), which contain activities that are recorded at a more granular level.
 The first group deals with communication, either Communication with customer or Communication with vendor.
@@
 7.1.4 Prediction performance for different timestamp encoding approaches
-Next to investigating strategies to embed activities of our TL-based PPM technique, we investigated five additional encoding strategies for the time-based feature duration since start time and compared them to our technique’s relative cross-domain encoding.
+Besides investigating strategies to embed activities of our TL-based PPM technique, we examined five additional encoding strategies for the time-based feature duration since start and compared them to our technique’s relative cross-domain encoding.
 These five encoding strategies fall into two categories: i) time-based encoding and ii) autoencoder-based encoding.
 The results in Table 8 indicate a clear trend towards the relative cross-domain encoding as it achieves the highest precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$ out of the six encoding strategies.
-The second-best encoding in all metrics is the autoencoder-based encoding.
+The encoding strategy resulting in the second-best prediction performance values in all metrics is the autoencoder-based encoding.
 Table 8: Prediction performance of our TL-based PPM technique with the proposed relative cross-domain encoding of the duration since start time feature and other encoding approaches for the intra-organizational use case (average and standard deviation over five runs).
@@
 7.1.5 Transfer relevance
-To provide an overview of how much target data should be available to train a model outperforming our transferred model (not fine-tuned), Figure 6 shows the $\text{AUC}_{\text{ROC}}$ for training a model solely on event data available in the target context with the first 1%, 2%, 5%, 10%, 20%, 50%, and 100% of cases using the all-mpnet-base-v2 (see Section 7.1.2) embedding model for activity encoding and the relative cross-domain approach (see Section 7.1.4) for timestamp encoding.
+To provide an overview of how much target data should be available to train a model outperforming our transferred model (not fine-tuned), Figure 6 shows the $\text{AUC}_{\text{ROC}}$ for training a model solely on event data available in the target context with the first 1%, 2%, 5%, 10%, 20%, 50%, and 100% of traces using the all-mpnet-base-v2 embedding model (see Section 7.1.2) for activity encoding and the relative cross-domain approach (see Section 7.1.4) for timestamp encoding.
 In terms of $\text{AUC}_{\text{ROC}}$, the transferred model achieves a higher $\text{AUC}_{\text{ROC}}$ compared to the majority of the models trained on any amount of available event data of the target context.
@@
 Similar to the intra-organizational use case, our proposed TL-based PPM technique outperforms all traditional PPM approaches in recall, F1 score, and $\text{AUC}_{\text{ROC}}$.
-The traditional PPM technique that employs the XGBoost model attains a higher recall, while the approach with the random forest model yields a superior MCC.
-Comparing the results of the transferred models with the baseline approach Train and test on $L_{T}$, transferred models come close to the prediction performance of the baseline approach Train and test on $L_{T}$ in most metrics.
+The traditional PPM technique that employs the XGBoost model attains a higher recall, whereas the approach with the random forest model yields a superior MCC.
+Comparing the results of the transferred models with the baseline approach Train and test on $L_{T}$, the transferred models come close to the prediction performance of the baseline approach Train and test on $L_{T}$ in most metrics.
 Table 9: Prediction performance of our proposed TL-based PPM technique and traditional PPM approaches for the inter-organizational use case (average and standard deviation over five runs).
@@
 7.2.3 Activity embedding quality
-Figure 7 shows the Euclidean distance between the activity embedding vectors from $L_{S}$ and $L_{T}$, created via the glove-wiki-gigaword-100 embedding model.
-[^26] The figures showing the Euclidean distance between activity embedding vectors created with the other embedding models of our setting can be found in the online repository.
-The source and target logs share only two common activities, Closed and RESOLVED, which are represented by an activity embedding distance of 0.00, as shown in Figure 5.
+Figure 7 shows the Euclidean distance between the embedding vectors of the activities from $L_{S}$ and $L_{T}$, created via the glove-wiki-gigaword-100 embedding model.
+[^26] The figures showing the Euclidean distance between activity embedding vectors created with the other embedding models of our setting can be found in the online repository of this work.
+The source and target logs share only two common activities, Closed and RESOLVED, which are represented by an activity embedding distance of 0.00.
 The other twelve activities in $L_{T}$ cannot be found in the source log $L_{S}$.
@@
 These are Create SW anomaly, Insert ticket, Require upgrade, Resolve SW anomaly, Resolve ticket, Schedule intervention, and Take in charge ticket.
-The other five activities in $L_{T}$, Assign seriousness, DUPLICATE, INVALID, VERIFIED, and Wait, are overall further apart from activities in $L_{S}$ based on the Euclidean distance between embedding vectors.
+The other five activities in $L_{T}$, Assign seriousness, DUPLICATE, INVALID, VERIFIED, and Wait, are overall further apart from activities in $L_{S}$ based on the Euclidean distance between the embedding vectors.
 However, the activities written in capital letters are outliers in terms of their frequency occurring in $L_{T}$ (see Table 5 in Section 6.1.2).
@@
 7.2.4 Prediction performance for different timestamp encoding approaches
-Similar to the intra-organizational use case, we also examined five additional encoding strategies for the time-based feature duration since start time, belonging to the categories time-based encoding and autoencoder-based encoding, and compared them to our technique’s relative cross-domain encoding.
+Similar to the intra-organizational use case, we also examined five additional encoding strategies for the time-based feature duration since start, belonging to the categories time-based encoding and autoencoder-based encoding, and compared them to our technique’s relative cross-domain encoding.
 The results in Table 8 indicate a clear trend towards the relative cross-domain encoding as it achieves the highest precision, recall, F1-score, MCC, and $\text{AUC}_{\text{ROC}}$ out of the six encoding strategies.
 The prediction performance of our technique using the time-based and autoencoder-based encoding strategies is comparably lower.
-Table 11: Prediction performance of our TL-based PPM technique with the proposed cross-domain encoding of the elapsed time attribute and other encoding approaches for the inter-organizational use case (average and standard deviation over five runs).
+Table 11: Prediction performance of our TL-based PPM technique with the proposed cross-domain encoding of the duration since start time feature and other encoding approaches for the inter-organizational use case (average and standard deviation over five runs).
 Encoding approach Precision Recall F1-score MCC $\text{AUC}_{\text{ROC}}$
@@
 7.2.5 Transfer relevance
-To provide an overview of how much event data of the target context should be available to train a model outperforming our transferred models (not fine-tuned), Figure 8 shows the $\text{AUC}_{\text{ROC}}$ for training a model solely on event data available in the target context with the first 1%, 2%, 5%, 10%, 20%, 50%, and 100% of cases using the glove-wiki-gigaword-100 (see Section 7.2.2) embedding model for activity encoding and the relative cross-domain approach (see Section 7.2.4) for time encoding.
+To provide an overview of how much event data of the target context should be available to train a model outperforming our transferred models (not fine-tuned), Figure 8 shows the $\text{AUC}_{\text{ROC}}$ for training a model solely on event data available in the target context with the first 1%, 2%, 5%, 10%, 20%, 50%, and 100% of traces using the glove-wiki-gigaword-100 embedding model (see Section 7.2.2) for activity encoding and the relative cross-domain approach (see Section 7.2.4) for time encoding.
 In terms of $\text{AUC}_{\text{ROC}}$, the transferred model achieves a higher prediction performance compared to the majority of the models trained on 1%, 2%, 5%, 10%, 20% of the available event data of the target context.
@@
 Second, our results in the inter- and intra-organizational use cases indicate that pre-trained embedding models can be a better choice to encode the activities for TL in PPM than using a sparse encoding technique, such as one-hot encoding.
-In addition, depending on the domain (event data) in the source and target contexts, different pre-trained embedding models lead to the best prediction performance of the transferred model in the target context.
+In addition, depending on the domain (event data) in the source and target context, different pre-trained embedding models lead to the best prediction performance of the transferred model in the target context.
 One possible explanation for this is that there is a correlation between the data used to pre-train the embedding models and the activity attribute values in the event data of the source and target contexts.
 Third, we show how embedding models can facilitate TL in PPM.
-By using these models, process activities from different business processes of the same type, having different names but similar semantic meanings, can be encoded accordingly.
+These models enable the encoding of process activities from different business processes of the same type that have different names but similar semantic meanings.
 This is possible with embedding models because they are trained in a self-supervised manner with auxiliary tasks (e.g., predicting a word in a sentence given the remaining words) to learn the representation of words while capturing their semantics.
@@
 Regarding the prediction target, a process-agnostic binary outcome prediction target, like the in-time prediction, is easier for the transfer than a process-specific outcome prediction or the next activity prediction.
-This is because the latter prediction targets are conceptually more different in the source and the target context.
-Concerning the considered process information, a transfer that relies solely on the control flow is typically easier than a transfer that relies on the control flow and additional time or data attributes.
+This is because the latter prediction targets differ more conceptually in the source and target contexts.
+Concerning the considered process information, a transfer that relies solely on the control flow is typically simpler than a transfer that relies on the control flow and additional time or data attributes.
 This is because the consideration of each additional attribute in the event data of the source and target contexts, with its values and distribution, increases the complexity.
@@
 Our work also has implications for practice.
-With a TL-based PPM technique like ours, the adoption of PPM may be fostered in practice.
+With a TL-based PPM technique like ours, the adoption of PPM can be encouraged in practice.
 This is specifically beneficial for organizations or areas in organizations without suitable resources, such as computational power, skills or training of employees, or event data for model training.
 These organizations can use the base model and further resources from another organization that already uses PPM successfully for similar business processes.
-This allows them to predict, for example, process outcomes or throughput times and, therefore, mitigate risks and decrease costs.
+This enables them to predict, for example, process outcomes or throughput times, and to use these predictions to mitigate risks and decrease costs.
 Furthermore, the proposed technique is scalable to incorporate the event data of multiple processes of the same type in the source and target contexts from a conceptual perspective.
@@
 However, it is conceivable that multiple event logs can be used as a domain in the source context for training our base model.
-Similarly, multiple event logs can be used as a domain in the target context to apply our base model.
+Similarly, multiple event logs can be used as different domains in the target context to which our base model is applied.
 To scale our technique in such a way, concepts from the field of federated learning can be incorporated (Verma et al.
 2019).
-As our technique uses an embedding model for activity encoding, scaling is easier than, for example, with a fixed encoding strategy like one-hot encoding, where each activity included in the event data of the source and target contexts has a specific position in the vector.
+As our technique uses an embedding model for activity encoding, scaling is easier than, for example, with a fixed encoding strategy like one-hot encoding, where each activity included in the event data of the source and target contexts has a specific position in the created feature vectors.
 8.3 Future research
@@
 First, fine-tuning of the embedding models used in our technique for feature-based TL can be investigated.
-In doing so, we suspect an improvement in the prediction performance of the model used in the target context.
+We suspect that this will improve the prediction performance of the model used in the target context.
 As the source context can comprise multiple business processes and their event data, certain strategies for fine-tuning can be realized to adapt the embedding models to multiple similar processes in the same context in a controlled way.
-For example, such a strategy could be to randomly take 1,000 traces of each business process’s event log to ensure a fair fine-tuning across multiple business processes with different amounts of event data.
+For example, such a strategy could be to randomly select 1,000 traces from each business process’s event log to ensure a fair fine-tuning across multiple business processes with varying amounts of event data.
 Second, the transfer of information about additional event log dimensions can be investigated.
@@
 2024) can be helpful to find a representation of multiple event log attributes, facilitating the process knowledge transfer from source to target contexts.
-Closely connected to the information used and its representation lies the question of which information is necessary for the knowledge transfer.
+The question of which information is necessary for the knowledge transfer is closely connected to the information used and its representation.
 This question can be investigated with explainable artificial intelligence (XAI) approaches, which are common in PPM (Stierle et al.
 2021).
-Third, the application of TL in the context of real business processes of different types can be investigated.
-In the case of those business processes, the complexity of TL is further increased as the business processes are more different from a structural and semantic perspective.
+Third, the application of TL in the context of different types of real business processes can be investigated.
+In the case of those business processes, the complexity of TL is further increased because they differ more structurally and semantically.
 For example, most of the activities, as well as additional data attributes in the event data of a business process in the source context, can differ from the event data of a business process in the target context.
@@
 As foundational models can support multiple modalities, types of process data going beyond text-transformed event data (e.g., process model collections or process-attached images, audio tracks, or video sequences) are additionally conceivable for model fine-tuning.
-To fine-tune a foundation model with event data from various organizations, the concept of federated learning can be essential to overcome data privacy restrictions (Verma et al.
+In order to fine-tune a foundation model with event data from various organizations, the concept of federated learning can be essential to overcome data privacy restrictions (Verma et al.
 2019).
 9 Conclusion
-In this paper, we proposed a TL-based technique for PPM, allowing organizations without suitable event data or other relevant resources to generate predictions in running business processes for effective decision support.
-For this purpose, we applied computational DSR (Rai et al.
+In this paper, we proposed a TL-based technique for PPM, allowing organizations without suitable event data or other relevant resources to create predictions in running business processes for effective decision support.
+For this purpose, we adopted the computational design science paradigm (Rai et al.
 2017) and designed an artifact in the form of a TL-based technique for PPM, consisting of three phases: i) initial model building on source, ii) transfer base model from source to target, and iii) online application of model to target.
-By instantiating our technique in the context of intra- and inter-organizational use cases, including event logs for ITSM processes from two organizations, we demonstrated and evaluated its prediction performance.
+Based on the instantiation of our technique in both a real-life intra- and an inter-organizational use case, including event logs for ITSM processes from two organizations, we evaluated its prediction performance.
 Our results suggest that knowledge of one business process can be transferred to a similar business process in the same or a different organization to enable effective PPM in the target context.
@@
 Table 12 shows that the complete model of our proposed TL-based technique clearly outperforms its sub-variants in terms of prediction performance when it is trained on $L_{S}$ and tested on $L_{T}$ for the inter-organizational use case.
-This confirms the importance of using both components in the prediction model for an effective process knowledge transfer from the source to the target contexts.
+This confirms the importance of using both components in the prediction model for an effective process knowledge transfer from source to target contexts.
 Table 12: Prediction performance of complete model compared to its sub-variants for the inter-organizational use case (average and standard deviation over five runs).
@@
 We also applied these three encoding approaches to two traditional PPM techniques, using LSTM and XGBoost models.
-We selected these two traditional PPM techniques because the first uses the same LSTM model as in our proposed technique but not in an TL-setting, and the second uses an XGBoost model, which performed relatively well in the first evaluation of our inter-case organizational use case (see Section 7.2.1.
+We selected these two traditional PPM techniques because the first uses the same LSTM model as in our proposed technique but not in an TL-setting, and the second uses an XGBoost model, which performed relatively well in the first evaluation of our inter-case organizational use case (see Section 7.2.1).
 Table 13 shows that our proposed technique with the index-based prefix encoding clearly outperforms all baseline approaches in terms of prediction performance on the test event data of the target context.
@@
 Table 14: Prediction performance for our TL-based PPM technique that encodes activity and timestamp information about a trace separately compared to a baseline technique that encodes activity and timestamp information about a trace jointly.
-The baseline technique was tested using the four pre-trained contextualized embeddings of our setup for the inter-organizational use case (average and standard deviation over five runs).
+The baseline technique was tested using the four pre-trained contextualized embedding models of our setup for the inter-organizational use case (average and standard deviation over five runs).
 Embedding model Precision Recall F1-score MCC $\text{AUC}_{\text{ROC}}$